### v2

# Judging the Judges: Evaluating Alignment and Vulnerabilities in LLMs-as-Judges

Abstract
Offering a promising solution to the scalability challenges associated with human evaluation, the LLM-as-a-judge paradigm is rapidly gaining traction as an approach to evaluating large language models (LLMs). However, there are still many open questions about the strengths and weaknesses of this paradigm, and what potential biases it may hold. In this paper, we present a comprehensive study of the performance of various LLMs acting as judges. [^1] Source code is available at https://github.com/UMass-Meta-LLM-Eval/llm_eval.git We leverage TriviaQA as a benchmark for assessing objective knowledge reasoning of LLMs and evaluate them alongside human annotations which we found to have a high inter-annotator agreement. Our study includes nine judge models and nine exam-taker models – both base and instruction-tuned. We assess the judge models’ alignment across different model sizes, families, and judge prompts. Among other results, our research rediscovers the importance of using Cohen’s kappa as a metric of alignment as opposed to simple percent agreement, showing that judges with high percent agreement can still assign vastly different scores. We find that both Llama-3 70B and GPT-4 Turbo have an excellent alignment with humans, but in terms of ranking exam taker models, they are outperformed by both JudgeLM-7B and the lexical matching method Contains, which have up to 34 points lower human alignment. Through error analysis and various other studies, including the effects of instruction length and leniency bias, we hope to provide useful lessons for using LLMs as judges in the future.
1 Introduction
Figure 1: (a) Scores assigned to all exam-taker models by the various judge models. (b) Average percent agreement (blue line) and Kappa scores (red bars) of judge models with human judges. Error bars annotate standard deviation across exam-taker models. Alignment is poor for most judge models, but both Llama3 70B and GPT-4 Turbo have Cohen’s Kappa coefficient that are indicative of excellent alignment (79 and 84, respectively), though still well below the human alignment score of 96.
Over the last few years, large language models (LLMs) have demonstrated remarkable capabilities across various domains (Radford et al., 2019; Brown et al., 2020; Achiam et al., 2023; AI@Meta, 2024, i.a.). As more and more new LLMs with different architectures and training methods continue to be released and their capabilities expand, accurately evaluating their performance and limitations becomes increasingly challenging (Zheng et al., 2024; Ohmer et al., 2024; Benchekroun et al., 2023; Madaan et al., 2024). The empirical evaluation of LLMs is particularly difficult due to the diversity of their outputs and the wide range of tasks they are used for (Zhang et al., 2024; Li et al., 2023a).
To evaluate LLMs, various methods have been proposed, that typically fall into one of two broad categories. Benchmarks such as MMLU (Hendrycks et al., 2021), TruthfulQA (Lin et al., 2021), and GSM8K (Cobbe et al., 2021) are used to evaluate specific capabilities of LLMs in an automated manner. Additionally, leaderboards like Chatbot Arena (Chiang et al., 2024) and Open LLM Leaderboard (Beeching et al., 2023) assign ranks to models considering pair-wise rankings of LLM outputs, done by humans or, in some cases, automated evaluation methods. Since both strategies involve evaluating free-form text responses generated by the LLMs, even in the first case, evaluating the responses is often just as challenging as generating them (see e.g. Chang et al., 2023).
One proposed solution to this problem is to use multiple-choice question (MCQ) benchmarks such as MMLU, and compare the log-probabilities of the potential answers rather than evaluating the generated answer directly. However, the MCQ paradigm ly limits the range of abilities that can be evaluated, and the setup increasingly diverges from how LLMs are used in practice. Alternatively, the use of lexical matching methods such as exact match (EM) or n-gram overlap to evaluate the responses are practical and cost-efficient approaches, but are susceptible to false negatives and often fail to adequately distinguish between responses with subtle differences that change their semantic meaning. This issue is exacerbated when evaluating instruction-tuned “chat” models that are fine-tuned to carry out conversations with humans in natural language, since their responses tend to be more verbose (Saito et al., 2023; Renze and Guven, 2024). For these reasons, human evaluation remains the gold standard for evaluating LLM responses. However, human evaluation is expensive, time-consuming, and often impractical in many use cases. As a result, it has increasingly become common practice to evaluate LLM responses using another LLM as a judge model (Lin et al., 2021; Islam et al., 2023; Chiang and Lee, 2023; Liusie et al., 2024). While there are promises of alignment between LLM judges and humans (Sottana et al., 2023; Zheng et al., 2024), there are also many open questions about the strengths and weaknesses of the paradigm.
In this work, we study the properties of LLMs as judges, comparing them with humans and automated evaluation methods. Contrary to prior work, we focus on a relatively ‘clean’ scenario in which inter-human agreement is as high as 96%, allowing us to distinguish ambiguity and subjectivity in the task itself from potential issues with the judge models. Using the knowledge benchmark TriviaQA (Joshi et al., 2017) as our playground, we investigate how nine different judge models with varying architectures and sizes judge nine different exam-taker models. Our main findings are:
  • Even in relatively straightforward setups, only the best models are suitable as judges. Out of the nine judge models we considered, only GPT-4 Turbo and Llama-3 70B showed very high alignment with humans, though even those judges’ alignment is still well behind the human alignment coefficient for the task (Figure 1).
  • Cohen’s kappa distinguishes judges better than percent alignment. In some cases, high percent agreement can still give very divergent scores (Figure 2).
  • Even Cohen’s kappa is not all telling. While GPT-4 Turbo and Llama-3 both have alignment scores that are considered excellent for discriminating different exam-taker models, their results are comparable to alternative cheaper approaches such as JudgeLM-7B and contains which have much lower alignment scores but more systematic biases (Figure 3)
Through detailed error analysis, we uncover additional insights into judge performance. Improved alignment appears to be driven by improved recall rates and reduced false negatives. However, judge models struggle with under-specified answers and tend to be lenient, affecting their evaluation consistency. They are also sensitive to the length and quality of prompts. And, surprisingly, even when the judge models are asked to evaluate an answer matching verbatim with a reference answer, many judge models still sometimes fail to evaluate it correctly.
Overall, our work showcases the strengths of the LLM-as-a-judge paradigm while also highlighting the need for caution against overreliance on alignment metrics, even in cases where they are high. Through error analysis, we also highlight several common failure cases that require attention. With this, we aim to contribute to a better general understanding of what is now becoming a mainstream paradigm for evaluating LLMs.
2 Related work
Various recent studies have used or considered using LLMs as judges for tasks such as evaluating story generation (Chiang and Lee, 2023), retrieval-augmented generation (Es et al., 2023), visual QA (Mañas et al., 2024), code comprehension (Zhiqiang et al., 2023), multilingual evaluation (Hada et al., 2023) and more general open-ended tasks (Zheng et al., 2024). Zhang et al. (2024) and Sottana et al. (2023) propose ways to standardise LLM evaluations and the role that judge models might play in such solutions. Several studies have demonstrated that state-of-the-art LLMs such as GPT-4 Turbo exhibit high alignment with human judgments (Sottana et al., 2023; Zheng et al., 2024), though others also illustrate that the paradigm is not yet without faults. Zeng et al. (2023) propose a benchmark for evaluating the performance of LLMs as judges, and other approaches have been proposed to improve LLM judges such that they are aligned well with humans (Shankar et al., 2024; Zhu et al., 2023).
Despite promising results in various settings, judge models still suffer from known issues of current LLMs such as hallucinations and factual errors (Ye et al., 2023; Turpin et al., 2023) and difficulty in following complex instructions (Li et al., 2023b; He et al., 2024). Furthermore, various studies have reported challenges such as position bias (Pezeshkpour and Hruschka, 2023; Zheng et al., 2023; Wang et al., 2023), verbosity bias (Saito et al., 2023) in their preferences, confusing evaluation criteria (Hu et al., 2024), or focusing more on the style and grammar compared to factuality (Wu and Aji, 2023). Recently, Liusie et al. (2024) have shown that LLMs perform better in comparative assessment compared to absolute scoring, which can be used for reliably measuring the relative performance of models (Liu et al., 2024) and creating classifiers for pairwise grading (Huang et al., 2024).
We follow up on this line of work and investigate the strengths and weaknesses of LLMs as judges. Unlike most prior work, we do not focus on pairwise comparisons of LLM outputs on open-ended tasks, but on comparisons of LLM outputs and reference answers. Since human alignment is high in this setting, this provides a clean playground to study the strengths and weaknesses of LLMs in detail. We also extend previous work by considering more LLMs, both as judges and LLMs to be evaluated.
Exam-taker models       (base & instruction-tuned) Llama-2 (7B, 13B and 70 B), Mistral 7B, GPT-4 Turbo
Judge models (instruction-tuned) Llama-2 (7B, 13B, and 70B), Llama-3 (8B and 70B),
Gemma 2B, Mistral 7B, JudgeLM 7B, GPT-4 Turbo
Table 1: The exam-taker models and judge models we use in our experiments. We consider a wide variety of judge models; to get a comprehensive overview of their (potential) biases, we consider exam-taker models of various sizes and types.
3 Methodology
To evaluate the strengths and weaknesses of the LLM-as-a-judge paradigm, we focus on a relatively controlled setup, in which judge models assess answers of exam-taker models on the knowledge benchmark TriviaQA (Joshi et al., 2017). We consider nine judge models and nine exam-taker models. More details about the dataset and models are provided in Appendix B.
Benchmark
As our testbed, we use the TriviaQA dataset (Joshi et al., 2017), consisting of 95K question-answer pairs sourced from 14 trivia and quiz league websites. Each question in the train and validation set is annotated with a list of short answers containing a minimal set of facts and evidence documents collected from Wikipedia and the Web. For our experiments, we use the validation set of the unfiltered partition of the benchmark, using the short answers as reference answers. We use the training set for few-shot examples. Since experiments require manual annotation of the exam-taker model responses, we use a random sample of $400$ questions from the dataset.
Exam-taker models
To understand the strengths and weaknesses of different judges, we benchmark pre-trained (base) and instruction-tuned (chat) exam-taker models across a wide variety of model sizes and examine the quality of the evaluations from different judge models. In particular, we consider Llama-2 (Touvron et al., 2023) in 7B, 13B, and 70B parameter sizes for both base and chat versions, Mistral 7B (Jiang et al., 2023) base and chat versions, and GPT-4 Turbo [^2] Accessed via the OpenAI API between Mar 19th, 2024 and May 19th, 2024. (Achiam et al., 2023) as the exam-taker models. The prompts for the exam-taker models contain five few-shot examples of (question, answer) pairs from the TriviaQA training set. The prompts for the instruction-tuned models additionally include a command signaling the model to answer the given question in a succinct manner similar to the provided examples. The prompts are provided in Appendix C.
Judge models
To get a comprehensive view of the strengths and weaknesses of judge models across different model sizes and architectures, we use instruction-tuned versions of Llama-2 (Touvron et al., 2023) in 7B, 13B, and 70B sizes, Llama-3 (AI@Meta, 2024) in 8B and 70B sizes, Mistral 7B (Jiang et al., 2023), GPT-4 Turbo (Achiam et al., 2023), Gemma 2B (Gemma Team et al., 2024), and JudgeLM 7B (Zhu et al., 2023) as judges. The judges are instructed to respond with only a single word, ‘‘correct’’ or ‘‘incorrect’’. The prompts can be found in Appendix D. An overview of al exam-taker models and judge models is shown in Table 1. For ease of reading, the judge models are depicted in a different font than the exam-taker models.
Baselines
As baselines, we use two commonly used lexical evaluation techniques – exact match (EM) and contains match (contains). For EM, a response is considered correct if the response exactly matches one of the reference answers for the given question. For contains, an answer is considered correct if at least one of the reference answers is a sub-string of the response string. Both EM and contains match are computed in a case-insensitive manner.
Alignment
We use two metrics to quantify alignment between judges: percent agreement and Cohen’s kappa coefficient (Cohen, 1960). Percent agreement expresses a simple percentage of the samples on which two annotators agree. Cohen’s kappa, denoted as $\mathbf{\kappa}$, is an alignment metric that also takes into account the possibility of chance agreement. It is generally considered to provide a more robust measure of alignment. Details about the computation of both metrics are given in Appendix E.
Human judgements
As a ground-truth assessment, we obtain human annotations for each exam-taker model answer. The inter-human alignment is calculated between three human judges using the answers to 600 questions by Llama-2 7B base, randomly sampled from the TriviaQA dataset (Joshi et al., 2017); the human guidelines can be found in Appendix F. We then determine collective “Human Judgment” through a majority vote. We found that the average alignment among human evaluators with the majority vote had a Cohen’s kappa score [^3] The values of the Cohen’s kappa coefficient have been scaled by 100×100$\times$100 × for easier comparison with percent agreement of $96.36\pm 1.67$, and the average percent agreement was $98.33\%\pm 0.76\%$. Given this near-perfect alignment score, we consider only one human evaluator per sample for the rest of our experiments, to reduce the overall cost of human annotations. The set of questions for which we obtain human annotations is identical for each exam-taker model.
4 Results
Figure 2: Delta evaluation score is calculated by taking judge score difference with human judgment; y-axes are in log scale. Fig (a) shows a skewed distribution for percent agreement and delta evaluation score. In Fig (b) we observe that highly aligned LLM judges with kappa > 0.8 exhibit low variability in scores. Conversely, judges with kappa < 0.8 demonstrate more variability, impacting their reliability.
In this section we discuss our main results, primarily focusing on the relationship between evaluations by various judge models and human evaluations (§ 4.1), and how that impacts their usability (§ 4.2). To do so, we evaluate their alignment with human judgment and assess how differently they rank the nine exam-taker models compared to humans. In Section 5, we further analyse their precision and recall to further investigate the types of errors that can be made by various judge models. Details about compute requirements and others costs for experiments are given in Appendix G.
4.1 Alignment between judge models and humans
We first compute $\kappa$ scores and percent agreement between the evaluations of each judge model and the human annotators. Figure 1 shows that while alignment is poor for most judge models, both Llama-3 70B and GPT-4 Turbo have $\mathbf{\kappa}$ scores – 79 and 84, respectively – that are considered to indicate excellent alignment. Nevertheless, there still is a significant disparity between human judgment and judge models: GPT-4 Turbo is 12 points behind human judgment. Notably, contains has a higher $\mathbf{\kappa}$ score than half of the judge models, while EM has the lowest alignment among all judges.
Cohen’s kappa vs percent agreement
In most cases, we observe that Cohen’s kappa and percent agreement are following the same trend, with the exception of the values for Gemma 2B and EM. Gemma 2B shows higher percent agreement compared to EM, yet it yields the lowest $\mathbf{\kappa}$ score within the ensemble. Furthermore, there is a significant difference in the actual values. For the percent agreement of judge models, we note a 30-point difference between human judgment and EM, while Cohen’s kappa exhibits a more substantial 53-point gap. This is also visible in the general decline of alignment scores: while Llama-3 8B has a $\kappa$ score of only 62, its percent agreement is still well above 80%. Overall, $\mathbf{\kappa}$ scores offer a more precise representation of diminishing trends in judge models compared to percent agreement.
Alignment vs assigned score
In Figure 2, we show the variation in scores assigned by the judge models to various exam-taker models for different values of percent agreement (Figure 2(a)) and Cohen’s kappa (Figure 2(b)). We can see that for $\mathbf{\kappa}$ > 80, the evaluation scores by judge models are close to the human evaluation scores for most of the judges, with a difference of up to 5 points in their assigned scores (complete results table provided in Appendix H). For percent agreement we observe deviations of up to 20 points in the evaluation scores for similar kappa alignment. Furthermore, we observe that the deviation from human-judgements can be quite distinct for different exam-taker models. In Figure 1(a), Gemma 2B, for instance, sometimes assigns higher scores than humans, and sometimes much lower. In the next section, we further explore this particular pattern.
4.2 Exploring systematic patterns in judge models
In the previous section, we have seen that none of the judge models we considered were aligned as well with humans as the humans themselves. Furthermore, as can be seen in Figure 2, the scores assigned by even the best aligned judge models can differ up to 10 points with the human-assigned scores. However, while this may limit – to some extent – the utility of using a judge models to get a perfect estimate of the exam-taker model’s capability on the benchmark, the judge models may still offer valuable insights to differentiate between different exam-taker models. If judges exhibit systematic biases such as – akin to a very strict teacher – consistently rating any exam-taker model lower, they will not assign identical scores but may assign identical rankings.
To evaluate this, we compare the rankings assigned by each judge model to the nine exam-taker models by computing their Spearman’s rank correlation coefficients $\rho$ (Spearman, 1904) with the human ranking. The rankings are shown in Figure 3(a), with $\rho$ values in Appendix J. These results show that contains demonstrates the highest alignment with the human ranking, swapping the ranks of only two out of nine models. Notably, contains performs on par with JudgeLM 7B (Zhu et al., 2023), a language model fine-tuned specifically for evaluating language model responses. They are closely followed by GPT-4 Turbo and Llama-3 70B, the judges with the best alignment. Remarkably, while GPT-4 Turbo and Llama-3 70B rank the same three models in positions three, four, and five as the human judge, they do so in different orders. Several other judges have rank correlations higher than 0.7; it appears they struggle to distinguish between poorer-performing exam-taker models, but do well at distinguishing between better-performing ones.
Figure 3: (a) Contains and JudgeLM retain 67% of the human-assigned ranking, closely followed by GPT-4 Turbo and LLama3-70B. All judges struggle to distinguish between the poor-performing exam-taker models; (b) False positives and negatives across different judge models in descending order of human alignment. Both false negatives and false positives generally increase as human alignment decreases, but well-aligned models tend to produce more false negatives than false positives.
5 Analysis
To better understand the judge models, we conduct multiple case studies aimed at identifying common errors and vulnerabilities.
5.1 Better aligned models: Recall gains, precision pains, and error spotlights
We first investigate the precision and recall of the judge models. We plot both – maintaining the ordering of Figure 1 – in Figure 4(a). It can be seen that the precision does not show any clear observable trend with increasing alignment, which can be further observed in Figure 3(b). The recall, on the other hand, shows an increasing trend, with more aligned models having comparatively fewer false negatives.
Figure 4: (a) Cohen’s kappa, precision, and recall for different judge models. Recall improves with increasing human alignment ($R^{2}$ = 0.98), though precision ($R^{2}$ = 0.0003) is not correlated with human alignment; (b) $\mathbf{\kappa}$ scores for each judge across different prompt templates. Except for GPT-4 Turbo, all judges struggle with too many detailed evaluation guidelines.
Next, we analyse the types of errors made by the judge models by manually annotating 900 outputs from Llama-7B Base with error codes, focusing on the top performers GPT-4 Turbo and Llama-3 70B. We then determine the percentage of each error type that are correctly judged to be incorrect by these two models. The results are shown in Table 2, where it can be observed that both GPT-4 Turbo and Llama-3 70B have a good error recall when the answers refer to an incorrect entity, or when too many entities are present. Under-specified and otherwise incorrect answers are most challenging for both judges, while answers with too few entities are judged relatively accurately by GPT-4 but less accurately by Llama-3 70B.
Error code Explanation Example Proportion GPT-4 recall Llama-3 70B recall
Incorrect entity
Response refers to a wrong entity
Henry VII, James I, Edward VI,
Mary I and Elizabeth I
86.9% 98.3% 96.6%
Under-specified
Response contains only part
of the answer
Henry VII, Henry VIII, Edward,
Mary, and Elizabeth
37.3% 33.9% 23.3%
Too few entities
Response contains too few entities
Henry VII, Edward VI,
Mary I and James I
2.47% 80.0% 60.0%
Too many entities
Response contains too many entities
Henry VII, Henry VIII, Edward VI,
Mary I, James I, and Elizabeth I
2.7% 90.1% 90.1%
Other
Response is incorrect but cannot
be put into any of the above buckets
I’m sorry but I do not know the
answer to that question
1.23% 20.0% 40.0%
Table 2: Error codes used to identify the types of errors made by exam-taker models when answering questions. The example question in this case is “Excluding Lady Jane Grey, who were the five monarchs of the House of Tudor?”, with the correct answer “Henry VII, Henry VIII, Edward VI, Mary I and Elizabeth I” (in any order).
5.2 Judge model sensitivity to prompt length and specificity
Next, we study the impact of the prompt on the ability of the judge models to perform an accurate assessment of exam-taker models, with two somewhat intertwined goals: 1) to study if the success of various judge models is related to the length of the prompt, and, 2) to study the degree to which the judgments of the judge models change with the specificity of the prompt.
We use four different prompt versions, varying in length and specificity. Each prompt instructs the judge models on how to evaluate responses, with complexity and detail increasing with the token count. The first two prompts (Without guidelines V1 and V2, with 45 and 58 tokens, respectively) simply ask to evaluate the responses, without any further information, while more elaborate guidance and examples are given in the longer prompts (Guidelines without examples and Guidelines with examples, with 245 and 301 tokens, respectively). All prompts are listed in Appendix K.
Figure 4(b) shows that Mistral 7B (Jiang et al., 2023), GPT-4 Turbo and Llama-3 70B exhibit relatively low variance in their agreement with humans as the level of information and the length of the prompt increases. For this task, top performers’ (GPT-4 Turbo and Llama-3 70B) implicit definition of a correct judgment seems well aligned with the provided instructions and thus shows high alignment with humans even if no specific instructions are provided. It can also be observed that only top performers appears to benefit from the more detailed instructions, with a slight upward trend, whereas the other models get less aligned with more instructions. This might be due to the less powerful judges not being able to follow many instructions in the prompt at the same time. Interestingly, in a follow-up experiment (see Appendix L), when the order of references provided to the judge models is shuffled, Figure 6(b) demonstrates that larger judge models consistently maintain their judgments regardless of the reference order, whereas smaller models except Mistral 7B are more sensitive to the reference order given in the prompt.
5.3 Evaluating controlled responses
Figure 5: We observe that judge models remain robust when exam-taker models produce responses identical to the prompt. However, this robustness diminishes in the presence of hallucinated responses such as "Yes" and "Sure" from the exam-taker models. Additionally, when assessing the gold standard answer, judges do not consistently arrive at the correct judgement 100% of the time.
Next, we perform some simple tests on the judge models by asking them to evaluate a set of dummy benchmark responses. For the first test, the answer to be evaluated for each question is one of the references from the dataset, verbatinm (the answer is thus always correct), while for the next three tests, the answer to be evaluated is always incorrect, with the dummy exam-taker model always responding with ‘‘Yes’’, and ‘‘Sure’’ for the second and the third tests, respectively, and simply repeating the question for the fourth test. Figure 5 shows that while some judge models are able to identify and correctly mark the answers as correct (for the first test) or incorrect (for the next three tests), some judges, notably Llama-2 70B incorrectly evaluates a significant number of dummy answers, even though it shows a relatively high alignment with humans on the benchmark evaluations (see Figure 1(b)).
We hypothesise that when the answers are plausible but incorrect (e.g. if the question asks about the name of the author of a book, and the exam-taker model gives the name of the wrong author), most judges are able to identify them as being incorrect (by comparing it with the reference answer). However, the judges might get confused about what they are supposed to evaluate if the answer is completely unrelated to the question (such as the words ‘‘Yes’’ and ‘‘Sure’’). It is possible that in this situation a judge model tries to evaluate one of the reference answers, thus marking it as correct, though further research is required to identify the cause of this behavior.
5.4 Leniency bias in judge models
Judge model $\kappa$ $P_{c}$ $P_{+}$
Gemma 2B 0.50 0.62 0.80
Llama-2 7B 0.66 0.68 0.36
Mistral 7B 0.72 0.75 0.75
Llama-2 13B 0.68 0.76 0.45
Llama-2 70B 0.80 0.79 0.94
Llama-3 8B 0.81 0.80 0.84
Llama-3 70B 0.84 0.84 0.79
GPT-4 Turbo 0.85 0.85 0.66
Figure 6: (a) Leniency bias across various judge models. To compute the leniency bias in judge models we estimate $P_{c}$ and $P_{+}$ for different judge models; (b) Consistency scores of each judge model for 3 random permutations of references in the prompt. The consistency score is the percentage of questions for which the judge model gives the same judgment for all 3 runs. We observe that the bigger models are more consistent as judges and less sensitive to reference orders in the prompt. More details are given in Appendix L.
To get a general sense of the inherent biases or misalignment in the evaluation criteria that might be present in the judge models, we estimate if they have a positive or negative bias in their judgment. To do so, we assume that a judge assigns the correct judgment (i.e. same evaluation as the ground truth) with a probability of $P_{c}$ and assigns the rest of the samples to be ‘‘correct’’ with a probability $P_{+}$, which we call their leniency bias. We estimate the values of $P_{c}$ and $P_{+}$ from the benchmark results,444 The theoretical derivation of the expressions for $P_{c}$ and $P_{+}$, as well as the empirical validation for their estimated values can be found in Appendix M. and show them in Figure 6(a). We observe that $P_{+}$ for most models is significantly higher than $0.5$, indicating a tendency of the judge models to evaluate responses as ‘‘correct’’ when their evaluation criteria are not completely aligned with the provided instructions.
6 Conclusion
In this work, we provide an extensive study of the properties of LLMs as judges, comparing them with human evaluation and automated evaluation methods. By focusing on a relatively ‘clean’ evaluation scenario in which inter-human agreement is high, we examine the potential issues with the LLM-as-a-judge paradigm separately from the ambiguity and subjectivity in the task itself. We find that even in relatively straightforward setups, smaller and more cost-efficient models are less effective judges compared to the best available LLMs – also if they are specifically trained as judges, such as JudgeLM – and that even the best LLMs fail to meet the consistency of humans. Furthermore, though previous work commonly used percent agreement, we found that Cohen’s kappa ($\mathbf{\kappa}$) distinguishes judges much better. However, we observe that even judges with excellent $\mathbf{\kappa}$ scores may require further scrutiny. While GPT-4 Turbo and Llama-3 both have excellent alignment scores, simpler and more cost-efficient approaches like JudgeLM and contains perform better when discriminating between the exam-taker models in terms of their ranking, despite having much lower alignment scores and more systematic biases. In further analyses, we find that: 1) LLMs tend to judge positively when in doubt, and this is more pronounced for small models than for larger ones; 2) Judge models with lower alignment lack precision rather than recall; 3) GPT-4 Turbo and Llama-3 70B are generally robust across different prompts, but are difficult to ‘steer’ in their judgments; 4) Some judge models can be easily fooled by dummy answers such as ‘Yes’ and ‘Sure’; 5) Judge models are better at detecting completely incorrect answers than partially incorrect ones.
Overall, this work adds to the realm of LLM evaluation research by assessing judges within a clearly defined and objective framework. Our results highlight the utility of using some LLMs as judges but also urge caution in blindly trusting their judgments, even if they are found to be well-aligned with humans. For practitioners using LLMs as judges – regardless of the setup – we recommend not only computing percent agreement, but also Cohen’s kappa, and pairing these with a qualitative analysis to ensure that conclusions from judge models are less susceptible to biases. We further elaborate on the limitations of our work in Appendix A. In the future, we plan to expand our work to increasingly more complex scenarios with more open-ended answers and variability, and more generally assess how consistent our findings are across dataset samples, benchmarks, and prompt templates.
Appendix A Limitations
In our work, we have shown potential pitfalls of using LLMs as judges, as well as the issues with commonly used metrics for quantifying their judgment quality. We have observed several trends and interesting phenomena regarding the behaviour of judge models in the LLM-as-a-judge paradigm, such as poor correlation or precision with alignment (§ 5.1), inability of smaller judge models to follow complex and nuanced instructions (§ 5.2), and tendencies of judge models to be lenient in their judgment when they are unsure (§ 5.4).
Broadly, our results show that caution and scrutiny are required in using LLMs as judges, but not all our specific conclusions about rankings may hold across different examples, different judge models and exam-taker model accuracies, and tasks. For example, on a small experiment with a different sample and different instructions, we observed that when we limited the questions to only those that have fewer than 10 or 20 references, there was a consistent 5-10 point drop in percent-alignment across all judge models, but the discriminative ability of the top-performing judges improved. In that example, in which the overall accuracy of the models was substantially lower than in the example in our main experiment, we also observed a closer match between Cohen’s kappa and percent alignment, suggesting that the accuracies of the exam-taker models being compared may play an important role in how well their scores match.
All in all, these differences underline how finicky using LLMs as judges can be, and with that confirm the overall conclusions of our study that much more work is needed to better understand the strengths and limitations of judge models across a wide range of scenarios and model accuracies. We consider assessing the strengths across multiple different samples and tasks, which would require many more human annotations, outside the scope of this paper and leave such experimentation for future work.
Appendix B Model and dataset details
In Table 3, we show the different models and datasets used in our experiments, along with version and license details.
Table 3: Version and license details for the different models and datasets used in experiments.
Asset Version License
TriviaQA mandarjoshi/trivia_qa apache-2.0
Llama-2 7B Base meta-llama/Llama-2-7b-hf llama2
Llama-2 7B Chat meta-llama/Llama-2-7b-chat-hf llama2
Llama-2 13B Base meta-llama/Llama-2-13b-hf llama2
Llama-2 13B Chat meta-llama/Llama-2-13b-chat-hf llama2
Llama-2 70B Base meta-llama/Llama-2-70b-hf llama2
Llama-2 70B Chat meta-llama/Llama-2-70b-chat-hf llama2
Mistral 7B Base mistralai/Mistral-7B-v0.1 apache-2.0
Mistral 7B Chat mistralai/Mistral-7B-Instruct-v0.2 apache-2.0
Llama-3 8B Chat meta-llama/Meta-Llama-3-8B-Instruct llama3
Llama-3 70B Chat meta-llama/Meta-Llama-3-70B-Instruct llama3
JudgeLM BAAI/JudgeLM-7B-v1.0 Non-commercial license
GPT-4 Turbo gpt-4-turbo-2024-04-09 N/A
Qwen 0.5B Chat Qwen/Qwen1.5-0.5B-Chat tongyi-qianwen
Qwen 1.8B Chat Qwen/Qwen1.5-1.8B-Chat tongyi-qianwen
Qwen 4B Chat Qwen/Qwen1.5-4B-Chat tongyi-qianwen
Qwen 7B Chat Qwen/Qwen1.5-7B-Chat tongyi-qianwen
Qwen 14B Chat Qwen/Qwen1.5-14B-Chat tongyi-qianwen
Qwen 72B Chat Qwen/Qwen1.5-72B-Chat tongyi-qianwen
Appendix C Model evaluation prompt templates
In Figure 7 and Figure 8, we show the prompt templates used for the base and chat exam-taker models during the question answering process.
Figure 7: Prompt template for base exam-taker models
Figure 8: Prompt template for Chat exam-taker models
Appendix D Judge LLM Prompt templates
In Figure 9, we show the prompt template used to guide the judge models during the evaluation process of a 400-question sample from the TriviaQA unfiltered dataset.
Figure 9: Prompt templates for the judge models
Appendix E Metrics for judge models
If one of the annotators is taken to be the reference, then the annotations of the other annotator can be categorized as true positives, false positives, true negatives, and false negatives, with the total number of each of them in a benchmark being represented by $T_{P},F_{P},T_{N},$ and $F_{N}$ respectively.
Percent agreement is simply the ratio of the numbers of times two annotators agree with each other relative to the total number of annotations. This ratio can have values between 00 and $1$. For the binary case, the alignment ratio $\rho$ is given as
$\rho=\frac{T_{P}+T_{N}}{T_{P}+F_{P}+T_{N}+F_{N}}.$ (1)
Cohen’s kappa coefficient, or Cohen’s kappa for short (Cohen, 1960), measures the alignment of two annotators while also taking into account the possibility of agreement by pure chance. This coefficient can have values between $-1$ and $1$, but is usually above 00 in most real-world situations. The value of Cohen’s kappa is given as
$\kappa$ $=\frac{p_{o}-p_{e}}{1-p_{e}}=\frac{2(T_{P}T_{N}-F_{P}F_{N})}{(T_{P}+F_{P})(T_{N}+F_{P})+(T_{P}+F_{N})(T_{N}+F_{N})}.$ (2)
Here, $p_{o}$ is the relative observed agreement, and $p_{e}$ is the hypothetical probability of chance agreement. This coefficient is considered to be a more robust measure of inter-annotator alignment, but also less interpretable in terms of what a particular value of $\kappa$ means. Generally, values of $\kappa$ in ranges $[0,0.2)$, $[0.2,0.4)$, $[0.4,0.6)$, $[0.6,0.8)$, and $[0.8,1)$ are considered to indicate no alignment, slight alignment, moderate alignment, substantial alignment, and near-perfect alignment respectively, with $\kappa=1$ indicating perfect alignment.
Appendix F Human Annotation Guidelines
We provide the guidelines used for human evaluation below.
Humn annotation guidelines You will be given a question, a set of reference answers and the answer given by an LLM. Your task is to judge if the answer given by the LLM is correct, as if you were the LLMs teacher grading their exam. An answer should be counted as correct if it is semantically equivalent to (one of the) reference answers. In doing so, please follow the following guidelines: Underspecified answers (e.g. "December" instead of "December 20") should be marked incorrect. Answers that have more information than requested (e.g. "December 20, in Paris" instead of "December 20") should be marked correct, provided the extra information is not incorrect or contrasting the rest of the answer. Answers with unnecessary verbosity but correct answers should be marked correct (E.g. ‘‘Thanks for asking this question! The correct answer is: ..."). If you have trouble judging whether the answer is correct, for instance because you feel you are lacking knowledge required to judge so, please indicate so by marking the answer "maybe correct" or ‘‘maybe incorrect", so that we can further review it.
Appendix G Experiment costs
The costs for the different experiments described in this work belong in three categories – GPU-hours for running open-source models on one or more Nvidia A100 GPUs, OpenAI credits for making API calls to OpenAI models, [^5] Pricing details for OpenAI models are available at https://openai.com/api/pricing/ and human hours for manual annotations of benchmark responses. The estimated costs for the final reported experiments are given in Table 4. In addition to this, previous unreported experiments and trials had an approximate cost of 120 GPU-hours, 100 USD in OpenAI credits, and 50 human hours, bringing the total experimental cost for this work to approximately 200 GPU-hours, USD 125 OpenAI credits, and 75 human annotation hours.
Table 4: Estimated costs for the final reported experiments. GPU-hours are in equivalent Nvidia A100 hours, OpenAI credits are in USD, and human hours are time spent in manual annotation.
Experiment GPU-hours OpenAI credits Human hours
Main benchmarks 5 2 -
Main evaluations 30 8 10
Human alignment 1 - 6
Error analysis 1.5 - 5
Controlled responses 15 - -
Leniency bias 5 5 -
Guideline bias 10 5 1
Reference bias 5 4 1
Total 72.5 24 23
Appendix H Judge Scores
We show the scores assigned by each judge model to each exam-taker model, visualised in Figure 1(a) in Table 5.
Table 5: Judge model score card for every exam-taker model.
Exam taker models
Llama2 Mistral GPT-4
Base Chat Base Instruct
Judge Models 7B 13B 70B 7B 13B 70B 7B
Llama 3 8B 68.75 76.00 85.25 52.91 75.0 78.5 62.5 66.00 76.25
Llama 3 70B 66.00 76.50 87.00 61.52 68.50 78.25 73.5 65.00 94.50
Llama 2 7B 61.75 71.25 80.25 58.48 77.25 58.00 68.25 80.50 86.75
Llama 2 13B 53.00 61.50 66.50 46.07 27.75 39.50 57.50 39.00 73.50
Llama 2 70B 68.25 76.30 86.50 64.05 80.04 82.25 77.00 74.25 95.25
Mistral 7B 67.75 75.00 85.25 64.55 68.75 77.50 74.50 68.50 92.75
Gemma 2B 67.00 74.00 84.00 100.00 44.25 75.25 100.00 79.50 99.75
JudgeLM 52.50 60.50 64.00 41.52 32.50 60.00 57.25 45.5 68.50
GPT-4 63.25 75.00 85.0 56.7 57.50 72.0 72.5 52.75 92.25
Exact Match 46.75 56.00 63.75 24.05 0.25 36.25 59.50 20.25 58.25
Contains Match 50.75 60.00 68.00 38.98 46.25 59.50 57.25 44.00 70.00
Human Eval 62.25 72.75 83.75 56.00 56.50 72.25 71.75 60.75 91.50
Appendix I Exam-taker model base vs chat analysis
Table 6: Scores of base and chat models by various judges
Judge models
Base-Chat pair EM Contains Human GPT-4 Turbo Llama-3 70B
Base Chat Base Chat Base Chat Base Chat Base Chat
Llama-2 7B 46.75 24.05 50.75 38.98 62.25 56.00 63.25 56.75 66.0 61.50
Mistral 7B 59.50 20.25 57.25 44.00 71.75 60.75 72.50 52.75 73.5 65.0
Llama-2 13B 56.00 0.25 60.00 46.25 72.75 56.50 75.00 57.50 76.50 68.50
Llama-2 70B 63.75 36.25 68.00 59.50 83.75 72.25 85.00 72.00 87.00 78.25
Given the human judgments we have available, we take the opportunity to investigate the performance differences between base and their corresponding chat models. In Table 6, we show the scores assigned by various judge models to four base-chat pairs. According to the default metric EM, the base models outperform the chat models by a large margin. Interestingly, while this difference gets smaller when the answers are judged by humans (second column) or GPT-4 Turbo, there is still a substantial difference for all four pairs, suggesting that the difference is not merely an effect of the increased verbosity of the chat models. Further evidence for that hypothesis is provided by Figure 10(b), in which we can see that while 14% of the errors are shared between the base-chat pairs, almost another 14% of the examples get judged correctly by the base models but not by the chat models, while the opposite happens in only 2.5% of the cases. We consider two alternative hypotheses:
  • i)
    The chat models have a worse understanding of the particular prompt format, which is tuned more to fit base models; or
  • ii)
    The chat models have ‘unlearned’ some knowledge during their alignment training.
To disentangle these two factors, we manually analyse 400 questions for Llama-2 70B and Llama-2 70B-chat, using our earlier error codes. The results, shown in Figure 10(a), sugest that, at least to some extent, the difference between base and chat models is in fact due to ‘unlearning’ of knowledge: while the number of errors is more or less equal among most categories, there is a stark difference in the incorrect entity category. Substantially more often than the base models, the chat models do answer the question with a semantically plausible but incorrect entity. In Table 7-9, we provide examples of such cases. The results do not show any evidence to support the first hypothesis: the number of errors where the answer cannot be parsed or is just entirely incorrect does not differ between base and chat models.
Figure 10: a) Distribution of incorrect question counts by error codes for the Llama2 70B Base vs Chat exam-taker models evaluated on 400 questions. b) Pie chart showing the percentage of questions categorized by the judgment from Base and Chat models.
Table 7: Knowledge unlearning example 1.
Question:
Which British artist’s works include ‘The First Real Target’?
References Peter Blake, Peter Balke, Sir Peter Blake
LLama-2 70B Base Peter Blake
LLama-2 70B Chat Patrick Caulfield
Mistral 7B Base David Hockney
Mistral 7B Chat Damien Hirst
Table 8: Knowledge unlearning example 2
Question:
Who was the first cricketer to score 10,000 test runs?
References Sunil Gavaskar, Sunil Manohar Gavaskar, SM Gavaskar, Sunny gavaskar, Gavaskar
LLama-2 70B Base Sunil Gavaskar
LLama-2 70B Chat Sachin Tendulkar
Mistral 7B Base Sachin Tendulkar
Mistral 7B Chat Sachin Tendulkar was the first cricketer to score 10,000 runs in Test matches.
Table 9: Knowledge unlearning example 3
Question:
‘Uncle Harry’s Coat’ was the first garment produced by which famous jacket manufacturer, based in Simonside, Newcastle Upon Tyne?
References Barbour
LLama-2 70B Base Barbour
LLama-2 70B Chat Barbour
Mistral 7B Base Barbour
Mistral 7B Chat Jack Walker & Sons
Appendix J Exam-taker model ranking correlation
In Table 10, we show the Spearman’s rank correlation coefficient (Spearman, 1904) ($\rho$) with human judgment. Since $\rho$ > 0.7 is considered well aligned, only Llama-7B and Gemma-2B have poor rank correlation with human judgment.
Table 10: Spearman Rank Correlation Coefficient $\rho$.
Judges $\rho$
Contains 0.98
JudgeLM-7B 0.98
GPT-4 0.93
Llama3-70B 0.93
Mistral-7B 0.92
Llama-13B 0.82
EM 0.78
Llama3-8B 0.77
Llama-70B 0.75
Llama-7B 0.39
Gemma-2B 0.21
Appendix K Too much info confuses judges
In Figure 11-14, we report the guidelines we used for the experiments in § 5.2. The simplest prompt used is Without Guidelines v1 (see Figure 11) where we define a sequential and structured process for the judge model. In Without Guidelines v2 (see Figure 12), we add an additional focus on the overall task and outcome as well. For Guidelines without examples (see Figure 13), we provide the judge models with detailed instructions about the task at hand, along with explicit guidelines on how to evaluate the answers. Additionally, for Guidelines with examples(see Figure 14), we also provide examples to the judge models for further reference.
Figure 11: Without Guidelines v1 prompt template for the judge models
Figure 12: Without Guidelines v2 prompt template for the judge models
Figure 13: Guidelines without examples Prompt template for the judge models
Figure 14: Guidelines with Examples Prompt template for the judge models
Appendix L Judge models are sensitive to reference order
We investigate the judges’ sensitivity to reference order by providing the same prompt, question and model response to the judge models, but shuffling the reference order in three different permutations. We compute the consistency score of the model as the percentage of questions for which it gives the same judgment all the 3 times. We observe that the model is more likely to evaluate an answer as correct if the corresponding reference appears early in the list of references (see Figure 15). The smaller judge models sometimes fail to capture all the information in the prompt, and provide judgement based on their own knowledge rather than going by the references (see Figure 16).
Figure 15: Example of Llama2-7B getting confused when the order of the references are changed
Figure 16: Example of Llama2-7B failing to identify the task by changing the order of the references.
Appendix M Leniency Bias
As described in § 5.4, for the purpose of the leniency bias experiments, we assume that a judge assigns the correct judgment with a probability of $P_{c}$ and randomly assigns the rest of the samples to be “correct” with a probability $P_{+}$. In this section, we derive the mathematical expressions for $P_{c}$ and $P_{+}$. We assume that in the case of misalignment between the evaluation criteria of guidelines and judge models, the probability of getting an evaluation of ‘‘correct’’ is independent of the actual correctness of the answer (i.e. the judge model effectively flips a coin to give out its judgement). For any given benchmark and judge model, we denote the ground-truth score as $s$, and the true positive and true negative rates as $t_{P}$ and $t_{N}$, respectively, all normalized to be between 00 and $1$.
Now, based on our assumptions, the true positives, where the exam-taker model response is correct, and also correctly identified by the judge model to be correct, would be comprised of two possible cases: 1) The judge evaluates it correctly according to the given evaluation criteria with a probability of $P_{c}$; and 2) The judge does not evaluate it according to the given criteria with a probability of $1-P_{c}$, but the evaluation still happens to be correct with a probability of $P_{+}$. With the total ratio of the correct responses being $s$, the true positive rate is therefore given by –
$t_{P}=s[P_{c}+(1-P_{c})P_{+}]$ (3)
Similarly, the true negatives, where the exam-taker model response is incorrect, and also correctly identified by the judge model to be incorrect, would also be comprised of two cases: 1) The judge evaluates it correctly according to the given evaluation criteria with a probability of $P_{c}$.2) The judge does not evaluate it according to the given criteria with a probability of $1-P_{c}$, but the evaluation still happens to be correct with a probability of $1-P_{+}$. With the total ratio of the incorrect responses being $1-s$, the true negative rate is therefore given by –
$t_{N}=(1-s)[P_{c}+(1-P_{c})(1-P_{+})].$ (4)
Using Equation 4, we can derive the following:
$t_{N}$ $=(1-s)[P_{c}+(1-P_{c})(1-P_{+})]$ (5)
$=P_{c}+1-P_{+}-P_{c}+P_{c}P_{+}-sP_{c}-s+sP_{+}+sP_{c}-sP_{c}P_{+}$ (6)
$=1-P_{+}+P_{c}P_{+}-s+sP_{+}-sP_{c}P_{+}$ (7)
$=1-s-P_{+}(1-P_{c}-s+sP_{c})$ (8)
$=1-s-P_{+}(1-s)(1-P_{c})$ (9)
$\implies P_{+}$ $=\frac{1-s-t_{N}}{(1-s)(1-P_{c})}$ (10)
$=\frac{1-\frac{t_{N}}{1-s}}{1-P_{c}}$ (11)
Substituting the value of $P_{+}$ in Equation 3, we get:
$t_{P}$ $=s[P_{c}+(1-P_{c})P_{+}]$ (12)
$=s\Bigg{[}P_{c}+(1-P_{c})\frac{1-\frac{t_{N}}{1-s}}{1-P_{c}}\Bigg{]}$ (13)
$=s\bigg{[}P_{c}+1-\frac{t_{N}}{1-s}\bigg{]}$ (14)
$\implies\frac{t_{P}}{s}$ $=P_{c}+1-\frac{t_{N}}{1-s}$ (15)
$\implies P_{c}$ $=\frac{t_{P}}{s}+\frac{t_{N}}{1-s}-1$ (16)
The values of $P_{c}$ and $P_{+}$ can be estimated from observed data using the derived expressions. In this experiment, we include Qwen models (Bai et al., 2023) of varying sizes, in our judge ensemble to increase the number of data points for this study. The estimated probabilities using this method, with human evaluation as the reference, are shown in Figure 17(a).
To validate these derived values, we observe the correlation between the estimated values of $P_{c}$ and Cohen’s kappa ($\kappa$). As shown in Figure 17(b), we observe that the estimated values of $P_{c}$ are highly correlated to the Cohen’s kappa values for the judge models, with a Pearson correlation coefficient of $0.98$.
Judge model $\kappa$ $P_{c}$ $P_{+}$
Qwen 1.8B 0.19 0.18 1.00
Qwen 0.5B 0.19 0.19 0.87
Gemma 7B 0.40 0.32 0.07
Gemma 2B 0.50 0.62 0.80
Qwen 4B 0.69 0.67 0.89
Llama-2 7B 0.66 0.68 0.36
Mistral 7B 0.72 0.75 0.75
Llama-2 13B 0.68 0.76 0.45
Llama-2 70B 0.80 0.79 0.94
Qwen 7B 0.80 0.79 0.77
Llama-3 8B 0.81 0.80 0.84
Qwen 72B 0.82 0.80 0.93
Qwen 14B 0.83 0.82 0.83
Llama-3 70B 0.84 0.84 0.79
GPT-4 Turbo 0.85 0.85 0.66
Figure 17: a) Estimated values of $P_{c}$ and $P_{+}$ for different judge models. b) Pearson’s correlation coefficient between $\kappa$ and $P_{c}$ for judge models.

### v3

@@
 However, there are still many open questions about the strengths and weaknesses of this paradigm, and what potential biases it may hold.
-In this paper, we present a comprehensive study of the performance of various LLMs acting as judges.
-[^1] Source code is available at https://github.com/UMass-Meta-LLM-Eval/llm_eval.git We leverage TriviaQA as a benchmark for assessing objective knowledge reasoning of LLMs and evaluate them alongside human annotations which we found to have a high inter-annotator agreement.
-Our study includes nine judge models and nine exam-taker models – both base and instruction-tuned.
-We assess the judge models’ alignment across different model sizes, families, and judge prompts.
-Among other results, our research rediscovers the importance of using Cohen’s kappa as a metric of alignment as opposed to simple percent agreement, showing that judges with high percent agreement can still assign vastly different scores.
-We find that both Llama-3 70B and GPT-4 Turbo have an excellent alignment with humans, but in terms of ranking exam taker models, they are outperformed by both JudgeLM-7B and the lexical matching method Contains, which have up to 34 points lower human alignment.
-Through error analysis and various other studies, including the effects of instruction length and leniency bias, we hope to provide useful lessons for using LLMs as judges in the future.
+In this paper, we present a comprehensive study of the performance of various LLMs acting as judges [^1] Source code is available at https://github.com/UMass-Meta-LLM-Eval/llm_eval Investigating thirteen judge models of different model sizes and families, judging answers of nine different ‘exam-taker models’ – both base and instruction-tuned – we find that only the best (and largest) models achieve reasonable alignment with humans.
+However, they are still quite far behind inter-human agreement and their assigned scores may still differ with up to 5 points from human-assigned scores.
+In terms of their ranking of the nine exam-taker models, instead, also smaller models and even the lexical metric contains may provide a reasonable signal.
+Through error analysis and other studies, we identify vulnerabilities in judge models, such as their sensitivity to prompt complexity and length, and a tendency toward leniency.
+The fact that even the best judges differ from humans in this comparatively simple setup suggest that caution may be wise when using judges in more complex setups.
+Lastly, our research rediscovers the importance of using alignment metrics beyond simple percent alignment, showing that judges with high percent agreement can still assign vastly different scores.
 1 Introduction
-Figure 1: (a) Scores assigned to all exam-taker models by the various judge models.
-(b) Average percent agreement (blue line) and Kappa scores (red bars) of judge models with human judges.
+Figure 1: Average scores assigned by judge models and alignment with human judges.
+(a) Scores assigned to all exam-taker models by the various judge models.
+(b) Average percent agreement (blue line) and Scott’s pi scores (red bars) of judge models with human judges (black line).
 Error bars annotate standard deviation across exam-taker models.
-Alignment is poor for most judge models, but both Llama3 70B and GPT-4 Turbo have Cohen’s Kappa coefficient that are indicative of excellent alignment (79 and 84, respectively), though still well below the human alignment score of 96.
+Llama3 70B, Llama3.1 70B and GPT-4 Turbo have Scott’s pi coefficient that are indicative of excellent alignment, but are still well below the human alignment score.
 Over the last few years, large language models (LLMs) have demonstrated remarkable capabilities across various domains (Radford et al., 2019; Brown et al., 2020; Achiam et al., 2023; AI@Meta, 2024, i.a.).
@@
 The empirical evaluation of LLMs is particularly difficult due to the diversity of their outputs and the wide range of tasks they are used for (Zhang et al., 2024; Li et al., 2023a).
-To evaluate LLMs, various methods have been proposed, that typically fall into one of two broad categories.
-Benchmarks such as MMLU (Hendrycks et al., 2021), TruthfulQA (Lin et al., 2021), and GSM8K (Cobbe et al., 2021) are used to evaluate specific capabilities of LLMs in an automated manner.
+To evaluate LLMs, various methods have been proposed, typically falling into one of two broad categories.
+First, benchmarks such as MMLU (Hendrycks et al., 2021), TruthfulQA (Lin et al., 2021), or GSM8K (Cobbe et al., 2021) are used to evaluate specific capabilities of LLMs in an automated manner.
 Additionally, leaderboards like Chatbot Arena (Chiang et al., 2024) and Open LLM Leaderboard (Beeching et al., 2023) assign ranks to models considering pair-wise rankings of LLM outputs, done by humans or, in some cases, automated evaluation methods.
 Since both strategies involve evaluating free-form text responses generated by the LLMs, even in the first case, evaluating the responses is often just as challenging as generating them (see e.g.
-Chang et al., 2023).
+Chang et al., 2023; Bavaresco et al., 2024).
 One proposed solution to this problem is to use multiple-choice question (MCQ) benchmarks such as MMLU, and compare the log-probabilities of the potential answers rather than evaluating the generated answer directly.
-However, the MCQ paradigm ly limits the range of abilities that can be evaluated, and the setup increasingly diverges from how LLMs are used in practice.
+However, the MCQ paradigm limits the range of abilities that can be evaluated, and the setup increasingly diverges from how LLMs are used in practice.
 Alternatively, the use of lexical matching methods such as exact match (EM) or n-gram overlap to evaluate the responses are practical and cost-efficient approaches, but are susceptible to false negatives and often fail to adequately distinguish between responses with subtle differences that change their semantic meaning.
-This issue is exacerbated when evaluating instruction-tuned “chat” models that are fine-tuned to carry out conversations with humans in natural language, since their responses tend to be more verbose (Saito et al., 2023; Renze and Guven, 2024).
+This issue is exacerbated when evaluating instruction-tuned “chat” models that are fine-tuned to carry out conversations with humans in natural language, since their responses tend to be more verbose (Saito et al., 2023; Renze & Guven, 2024).
 For these reasons, human evaluation remains the gold standard for evaluating LLM responses.
 However, human evaluation is expensive, time-consuming, and often impractical in many use cases.
-As a result, it has increasingly become common practice to evaluate LLM responses using another LLM as a judge model (Lin et al., 2021; Islam et al., 2023; Chiang and Lee, 2023; Liusie et al., 2024).
+As a result, it has increasingly become common practice to evaluate LLM responses using another LLM as a judge model (Lin et al., 2021; Islam et al., 2023; Chiang & Lee, 2023; Liusie et al., 2024).
 While there are promises of alignment between LLM judges and humans (Sottana et al., 2023; Zheng et al., 2024), there are also many open questions about the strengths and weaknesses of the paradigm.
 In this work, we study the properties of LLMs as judges, comparing them with humans and automated evaluation methods.
-Contrary to prior work, we focus on a relatively ‘clean’ scenario in which inter-human agreement is as high as 96%, allowing us to distinguish ambiguity and subjectivity in the task itself from potential issues with the judge models.
-Using the knowledge benchmark TriviaQA (Joshi et al., 2017) as our playground, we investigate how nine different judge models with varying architectures and sizes judge nine different exam-taker models.
+Contrary to prior work, we focus on a clean scenario in which human alignment is very high, allowing us to distinguish ambiguity and subjectivity in the task itself from potential issues with the judge models.
+Using the knowledge benchmark TriviaQA (Joshi et al., 2017) as our playground, we investigate how thirteen different judge models with varying architectures and sizes judge nine different exam-taker models.
 Our main findings are:
-  • Even in relatively straightforward setups, only the best models are suitable as judges.
-Out of the nine judge models we considered, only GPT-4 Turbo and Llama-3 70B showed very high alignment with humans, though even those judges’ alignment is still well behind the human alignment coefficient for the task (Figure 1).
-  • Cohen’s kappa distinguishes judges better than percent alignment. In some cases, high percent agreement can still give very divergent scores (Figure 2).
-  • Even Cohen’s kappa is not all telling.
-While GPT-4 Turbo and Llama-3 both have alignment scores that are considered excellent for discriminating different exam-taker models, their results are comparable to alternative cheaper approaches such as JudgeLM-7B and contains which have much lower alignment scores but more systematic biases (Figure 3)
-Through detailed error analysis, we uncover additional insights into judge performance.
+  • Even in clean and straightforward setups, only the best models have high alignment scores.
+Out of the thirteen judge models we considered, only GPT-4 Turbo, Llama-3.1 70B and Llama-3 70B showed very high alignment with humans.
+Also for those judges, though, alignment is still well behind the human alignment coefficient for the task (Figure 1).
+  • Scott’s $\mathbf{\pi}$ distinguishes judges better than percent alignment.
+In terms of percent alignment, judges are rarely discriminable, while Scott’s $\mathbf{\pi}$ provides a more informative signal.
+In some cases, high percent agreement can still give scores that differ 10-20 points from the human-assigned scores (Figure 2).
+  • Also Scott’s $\mathbf{\pi}$ is not all telling.
+While GPT-4 Turbo and Llama-3 both have alignment scores that are considered excellent, their scores still differ up to 5 points from human-assigned scores.
+Furthermore, when it comes to discriminating different exam-taker models, their results are comparable to alternative cheaper approaches such as Mistral 7B and contains, which have much lower alignment scores but more consistent biases (Figure 3).
+Through detailed analysis (§ 5), we uncover additional insights into judge performance.
 Improved alignment appears to be driven by improved recall rates and reduced false negatives.
@@
 2 Related work
-Various recent studies have used or considered using LLMs as judges for tasks such as evaluating story generation (Chiang and Lee, 2023), retrieval-augmented generation (Es et al., 2023), visual QA (Mañas et al., 2024), code comprehension (Zhiqiang et al., 2023), multilingual evaluation (Hada et al., 2023) and more general open-ended tasks (Zheng et al., 2024).
+Various recent studies have used or considered using LLMs as judges for tasks such as evaluating story generation (Chiang & Lee, 2023), retrieval-augmented generation (Es et al., 2023), visual QA (Mañas et al., 2024), code comprehension (Zhiqiang et al., 2023), multilingual evaluation (Hada et al., 2023) and more general open-ended tasks (Zheng et al., 2024).
 Zhang et al.
@@
 Despite promising results in various settings, judge models still suffer from known issues of current LLMs such as hallucinations and factual errors (Ye et al., 2023; Turpin et al., 2023) and difficulty in following complex instructions (Li et al., 2023b; He et al., 2024).
-Furthermore, various studies have reported challenges such as position bias (Pezeshkpour and Hruschka, 2023; Zheng et al., 2023; Wang et al., 2023), verbosity bias (Saito et al., 2023) in their preferences, confusing evaluation criteria (Hu et al., 2024), or focusing more on the style and grammar compared to factuality (Wu and Aji, 2023).
+Furthermore, various studies have reported challenges such as position bias (Pezeshkpour & Hruschka, 2023; Zheng et al., 2023; Wang et al., 2023), verbosity bias (Saito et al., 2023) in their preferences, confusing evaluation criteria (Hu et al., 2024), or focusing more on the style and grammar compared to factuality (Wu & Aji, 2023).
 Recently, Liusie et al.
@@
 We also extend previous work by considering more LLMs, both as judges and LLMs to be evaluated.
+Table 1: Exam-taker models and judge models We consider a wide variety of exam-taker models and judge models; to get a in-depth overview of their abilities, we consider exam-taker models of various sizes & types.
 Exam-taker models       (base & instruction-tuned) Llama-2 (7B, 13B and 70 B), Mistral 7B, GPT-4 Turbo
-Judge models (instruction-tuned) Llama-2 (7B, 13B, and 70B), Llama-3 (8B and 70B),
+Judge models (instruction-tuned) Llama-2 (7B, 13B, 70B),
+Llama-3 (8B, 70B), Llama-3.1 (8B, 70B),
 Gemma 2B, Mistral 7B, JudgeLM 7B, GPT-4 Turbo
-Table 1: The exam-taker models and judge models we use in our experiments.
-We consider a wide variety of judge models; to get a comprehensive overview of their (potential) biases, we consider exam-taker models of various sizes and types.
+Judge models (lexical) Exact Match (EM), Contains
 3 Methodology
-To evaluate the strengths and weaknesses of the LLM-as-a-judge paradigm, we focus on a relatively controlled setup, in which judge models assess answers of exam-taker models on the knowledge benchmark TriviaQA (Joshi et al., 2017).
-We consider nine judge models and nine exam-taker models.
-More details about the dataset and models are provided in Appendix B.
-Benchmark
+To evaluate the strengths and weaknesses of the LLM-as-a-judge paradigm, we focus on a comparatively controlled setup, in which judge models assess answers of exam-taker models on the knowledge benchmark TriviaQA (Joshi et al., 2017).
+With this methodological design, it is possible to focus on the abilities of the judges in isolation, without having to address human disagreement and error at the same time.
+In this section, we elaborate the main aspects of our methodology.
+Evaluation data
 As our testbed, we use the TriviaQA dataset (Joshi et al., 2017), consisting of 95K question-answer pairs sourced from 14 trivia and quiz league websites.
@@
 Since experiments require manual annotation of the exam-taker model responses, we use a random sample of $400$ questions from the dataset.
+In Appendix I, we show with a bootstrapping test that this sample size has low variance for our main result.
+Through experiments described in § 3, we establish that humans have high agreement on judgements of answers given to the questions in the benchmark.
 Exam-taker models
-To understand the strengths and weaknesses of different judges, we benchmark pre-trained (base) and instruction-tuned (chat) exam-taker models across a wide variety of model sizes and examine the quality of the evaluations from different judge models.
-In particular, we consider Llama-2 (Touvron et al., 2023) in 7B, 13B, and 70B parameter sizes for both base and chat versions, Mistral 7B (Jiang et al., 2023) base and chat versions, and GPT-4 Turbo [^2] Accessed via the OpenAI API between Mar 19th, 2024 and May 19th, 2024.
+To understand the strengths and weaknesses of different judges, we consider answers of pre-trained (base) and instruction-tuned (chat) ‘exam-taker models’ across a wide variety of model sizes.
+In particular, we consider Llama-2 (Touvron et al., 2023) in 7B, 13B, and 70B parameter sizes for both base and chat versions, Mistral 7B (Jiang et al., 2023) base and chat versions, and GPT-4 Turbo [^2] Accessed via the OpenAI API between Mar 19th, 2024 and Sep 20, 2024.
 (Achiam et al., 2023) as the exam-taker models.
@@
 The prompts for the instruction-tuned models additionally include a command signaling the model to answer the given question in a succinct manner similar to the provided examples.
-The prompts are provided in Appendix C.
+The prompts are provided in Appendix D.
 Judge models
-To get a comprehensive view of the strengths and weaknesses of judge models across different model sizes and architectures, we use instruction-tuned versions of Llama-2 (Touvron et al., 2023) in 7B, 13B, and 70B sizes, Llama-3 (AI@Meta, 2024) in 8B and 70B sizes, Mistral 7B (Jiang et al., 2023), GPT-4 Turbo (Achiam et al., 2023), Gemma 2B (Gemma Team et al., 2024), and JudgeLM 7B (Zhu et al., 2023) as judges.
+To get a comprehensive view of the strengths and weaknesses of judge models across different model sizes and architectures, we use instruction-tuned versions of Llama-2 (Touvron et al., 2023) in 7B, 13B, and 70B sizes, Llama-3 (AI@Meta, 2024) in 8B and 70B sizes, Llama-3.1 (Dubey et al., 2024) in 8B and 70B sizes, Mistral 7B (Jiang et al., 2023), GPT-4 Turbo (Achiam et al., 2023), Gemma 2B (Gemma Team et al., 2024), and JudgeLM 7B (Zhu et al., 2023) as judges.
+To maintain parity with human and judge evaluation, judge prompts were built from human guidelines in Appendix G.
 The judges are instructed to respond with only a single word, ‘‘correct’’ or ‘‘incorrect’’.
-The prompts can be found in Appendix D.
-An overview of al exam-taker models and judge models is shown in Table 1.
+An overview of all exam-taker models and judge models is shown in Table 1.
 For ease of reading, the judge models are depicted in a different font than the exam-taker models.
@@
 Alignment
-We use two metrics to quantify alignment between judges: percent agreement and Cohen’s kappa coefficient (Cohen, 1960).
+We use two metrics to quantify alignment between judges: percent agreement and Scott’s Pi coefficient (Scott, 1955).
+[^3] In an earlier version of this paper, we used Cohen’s kappa (Cohen, 1960) to measure alignment.
+It has since come to our attention that – despite it’s widespread use – this metric has some well-documented theoretical issues (e.g.
+Pontius & Millones, 2011; Chicco et al., 2021).
+For the interested reader, we elaborate on these issues in Appendix B.
 Percent agreement expresses a simple percentage of the samples on which two annotators agree.
-Cohen’s kappa, denoted as $\mathbf{\kappa}$, is an alignment metric that also takes into account the possibility of chance agreement.
-It is generally considered to provide a more robust measure of alignment.
-Details about the computation of both metrics are given in Appendix E.
+Scott’s Pi, denoted as Scott’s $\mathbf{\pi}$, is an alignment metric that corrects for chance agreement between two annotators and is considered to provide a more robust measure of alignment.
+Details about the computation of both metrics are given in Appendix F.
 Human judgements
 As a ground-truth assessment, we obtain human annotations for each exam-taker model answer.
-The inter-human alignment is calculated between three human judges using the answers to 600 questions by Llama-2 7B base, randomly sampled from the TriviaQA dataset (Joshi et al., 2017); the human guidelines can be found in Appendix F.
+The inter-human alignment is calculated between three human judges using the answers to 1200 randomly sampled questions answers; the human guidelines can be found in Appendix G.
 We then determine collective “Human Judgment” through a majority vote.
-We found that the average alignment among human evaluators with the majority vote had a Cohen’s kappa score [^3] The values of the Cohen’s kappa coefficient have been scaled by 100×100$\times$100 × for easier comparison with percent agreement of $96.36\pm 1.67$, and the average percent agreement was $98.33\%\pm 0.76\%$.
+The average alignment among human evaluators with the majority vote had a Scott’s $\mathbf{\pi}$ of $96.2\pm 1.07$,444coefficient scaled by $100$ for easier comparison with percent alignment.
+and the average percent agreement was $98.52\%\pm 0.42\%$.
+The details of this experiment are mentioned in Appendix A.
 Given this near-perfect alignment score, we consider only one human evaluator per sample for the rest of our experiments, to reduce the overall cost of human annotations.
@@
 4 Results
-Figure 2: Delta evaluation score is calculated by taking judge score difference with human judgment; y-axes are in log scale.
-Fig (a) shows a skewed distribution for percent agreement and delta evaluation score.
-In Fig (b) we observe that highly aligned LLM judges with kappa > 0.8 exhibit low variability in scores.
-Conversely, judges with kappa < 0.8 demonstrate more variability, impacting their reliability.
+Figure 2: Difference with human evaluation scores versus alignment metric.
+The delta evaluation score is the difference between the judge and the human score; y-axes are in log scale.
+Percent alignment (left) shows a very skewwed distribution, making it difficult to distinguish models.
+Scott’s $\mathbf{\pi}$ (left) provides a clearer difference between models, and is more indicative of deviation of the gold score.
 In this section we discuss our main results, primarily focusing on the relationship between evaluations by various judge models and human evaluations (§ 4.1), and how that impacts their usability (§ 4.2).
@@
 In Section 5, we further analyse their precision and recall to further investigate the types of errors that can be made by various judge models.
-Details about compute requirements and others costs for experiments are given in Appendix G.
+Details about compute requirements and others costs for experiments are given in Appendix H.
 4.1 Alignment between judge models and humans
-We first compute $\kappa$ scores and percent agreement between the evaluations of each judge model and the human annotators.
-Figure 1 shows that while alignment is poor for most judge models, both Llama-3 70B and GPT-4 Turbo have $\mathbf{\kappa}$ scores – 79 and 84, respectively – that are considered to indicate excellent alignment.
-Nevertheless, there still is a significant disparity between human judgment and judge models: GPT-4 Turbo is 12 points behind human judgment.
-Notably, contains has a higher $\mathbf{\kappa}$ score than half of the judge models, while EM has the lowest alignment among all judges.
-Cohen’s kappa vs percent agreement
-In most cases, we observe that Cohen’s kappa and percent agreement are following the same trend, with the exception of the values for Gemma 2B and EM.
-Gemma 2B shows higher percent agreement compared to EM, yet it yields the lowest $\mathbf{\kappa}$ score within the ensemble.
-Furthermore, there is a significant difference in the actual values.
-For the percent agreement of judge models, we note a 30-point difference between human judgment and EM, while Cohen’s kappa exhibits a more substantial 53-point gap.
-This is also visible in the general decline of alignment scores: while Llama-3 8B has a $\kappa$ score of only 62, its percent agreement is still well above 80%.
-Overall, $\mathbf{\kappa}$ scores offer a more precise representation of diminishing trends in judge models compared to percent agreement.
-Alignment vs assigned score
-In Figure 2, we show the variation in scores assigned by the judge models to various exam-taker models for different values of percent agreement (Figure 2(a)) and Cohen’s kappa (Figure 2(b)).
-We can see that for $\mathbf{\kappa}$ > 80, the evaluation scores by judge models are close to the human evaluation scores for most of the judges, with a difference of up to 5 points in their assigned scores (complete results table provided in Appendix H).
-For percent agreement we observe deviations of up to 20 points in the evaluation scores for similar kappa alignment.
-Furthermore, we observe that the deviation from human-judgements can be quite distinct for different exam-taker models.
+We start by computing Scott’s $\mathbf{\pi}$ scores and percent agreement between the evaluations of each judge model and the human annotators.
+We show the result in Figure 1.
+We observe that percent alignment is high for virtually all models, with the exception of Gemma 2B and EM.
+Scott’s $\mathbf{\pi}$, on the other hand, has low values for most models, though its value is in the high 80s for Llama-3 70B, Llama-3.1 70B and GPT-4 Turbo.
+Nevertheless, there still is a significant disparity between human judgment and judge models: the best scoring judge, Llama-3 70B, is 8 points behind human judgment.
+Notably, EM has the most variance in alignment, while Gemma 2B has the lowest alignment amongst all judges.
+In most cases, we observe that Scott’s $\mathbf{\pi}$ and percent agreement are following the same trend, with the exception of the values for Gemma 2B and EM.
+Gemma 2B shows higher percent agreement compared to EM, yet it yields the lowest Scott’s $\mathbf{\pi}$ score within the ensemble.
+For the percent agreement of judge models, we note a 26-point difference between human judgment and EM, while Scott’s $\mathbf{\pi}$ exhibits a more substantial 64-point gap.
+This is also visible in the general decline of alignment scores: while Llama-3 8B has a Scott’s $\mathbf{\pi}$ score of only 59, its percent agreement is still well above 80%.
+Overall, Scott’s $\mathbf{\pi}$ appears to be better able of discriminating various judge models, showing more divergence across the tested judges.
+To understand how indicative the two alignment metrics are of the expected accuracy of the overall judgement of the models, we plot, for each judge model and exam-taker model, the difference between the score assigned by the judge and the score assigned by a human.
+In the figure, we can see that for Scott’s $\mathbf{\pi}$ values higher than 80, the evaluation scores are comparatively close to the human evaluation scores, with a difference of up to 5 points in their assigned scores (complete results table provided in Appendix J).
+For percent alignment, on the other hand, even judges that have more than 90% may still differ more than 10 points in their assigned score.
+Interestingly, the deviation from human-judgements for a single judge model can be quite different depending on the exam-taker model.
 In Figure 1(a), Gemma 2B, for instance, sometimes assigns higher scores than humans, and sometimes much lower.
 In the next section, we further explore this particular pattern.
-4.2 Exploring systematic patterns in judge models
-In the previous section, we have seen that none of the judge models we considered were aligned as well with humans as the humans themselves.
-Furthermore, as can be seen in Figure 2, the scores assigned by even the best aligned judge models can differ up to 10 points with the human-assigned scores.
+4.2 Exploring consistent patterns in judge models
+In the previous section, we have seen that none of the judge models we considered were aligned with humans as well as the humans were aligned amongst themselves.
+Furthermore, as can be seen in Figure 2, the scores assigned by even the best aligned judge models can differ up to 5 points with the human-assigned scores.
 However, while this may limit – to some extent – the utility of using a judge models to get a perfect estimate of the exam-taker model’s capability on the benchmark, the judge models may still offer valuable insights to differentiate between different exam-taker models.
-If judges exhibit systematic biases such as – akin to a very strict teacher – consistently rating any exam-taker model lower, they will not assign identical scores but may assign identical rankings.
+If judges exhibit consistent biases such as – akin to a very strict teacher – consistently rating any exam-taker model lower, they will not assign identical scores but may assign identical rankings.
 To evaluate this, we compare the rankings assigned by each judge model to the nine exam-taker models by computing their Spearman’s rank correlation coefficients $\rho$ (Spearman, 1904) with the human ranking.
-The rankings are shown in Figure 3(a), with $\rho$ values in Appendix J.
-These results show that contains demonstrates the highest alignment with the human ranking, swapping the ranks of only two out of nine models.
-Notably, contains performs on par with JudgeLM 7B (Zhu et al., 2023), a language model fine-tuned specifically for evaluating language model responses.
-They are closely followed by GPT-4 Turbo and Llama-3 70B, the judges with the best alignment.
-Remarkably, while GPT-4 Turbo and Llama-3 70B rank the same three models in positions three, four, and five as the human judge, they do so in different orders.
-Several other judges have rank correlations higher than 0.7; it appears they struggle to distinguish between poorer-performing exam-taker models, but do well at distinguishing between better-performing ones.
-Figure 3: (a) Contains and JudgeLM retain 67% of the human-assigned ranking, closely followed by GPT-4 Turbo and LLama3-70B.
-All judges struggle to distinguish between the poor-performing exam-taker models; (b) False positives and negatives across different judge models in descending order of human alignment.
-Both false negatives and false positives generally increase as human alignment decreases, but well-aligned models tend to produce more false negatives than false positives.
+We show the rankings in Figure 3(a), with $\rho$ and corresponding $\sigma$ values in Appendix L.
+Most judge models have rank correlations higher than 0.7; it appears they struggle to distinguish between poorer-performing exam-taker models, but do well at distinguishing between better-performing ones.
+Notably, the results show that several models that assign scores quite divergent from humans and have poor alignment on the sample level are very aligned in terms of the rankings they assign.
+Specifically, both contains and Mistral 7B, with Scott’s $\mathbf{\pi}$ values of 64 and 66, respectively, exhibit very high rank correlation with the human scores ($\rho$ 0.99 and 0.98, respectively, with $\sigma$ 0.02 and 0.03).
+With that, these judges perform on par with GPT-4 Turbo and outperform the better Llama judges – though with lower significance values – indicating that identifying which models are better should not be equated to assigning them the correct score.
+Figure 3: Judge rankings and true/false positives and negatives.
+(a) Assigned exam-taker model rankings assigned by highly human aligned judges.
+Contains stays closely to human-assigned rankings, as well as GPT-4 Turbo and Mistral 7B.
+(b) False positives and negatives across different judge models, in descending order of human alignment.
+Both false negatives and false positives increase as human alignment decreases, but well-aligned models tend to produce more false negatives than false positives.
 5 Analysis
-To better understand the judge models, we conduct multiple case studies aimed at identifying common errors and vulnerabilities.
-5.1 Better aligned models: Recall gains, precision pains, and error spotlights
+To better understand the judge models, we conduct multiple case studies aimed at identifying common errors and vulnerabilities in the judges we investigate.
+Specifically, we study their precision and recall and error types (§ 5.1), their sensitivity to the instruction prompt prompt (§ 5.2), how they respond to controlled resposes of specific types (§ 5.3), and the extent to which they have a leniency bias (§ 5.4).
+5.1 Better aligned models: Precision and recall gains with error spotlights
 We first investigate the precision and recall of the judge models.
-We plot both – maintaining the ordering of Figure 1 – in Figure 4(a).
-It can be seen that the precision does not show any clear observable trend with increasing alignment, which can be further observed in Figure 3(b).
-The recall, on the other hand, shows an increasing trend, with more aligned models having comparatively fewer false negatives.
-Figure 4: (a) Cohen’s kappa, precision, and recall for different judge models.
-Recall improves with increasing human alignment ($R^{2}$ = 0.98), though precision ($R^{2}$ = 0.0003) is not correlated with human alignment; (b) $\mathbf{\kappa}$ scores for each judge across different prompt templates.
-Except for GPT-4 Turbo, all judges struggle with too many detailed evaluation guidelines.
+Maintaining the ordering of Figure 1, we plot both in Figure 4(a).
+We can see that both precision and recall exhibit a moderate increasing trend as alignment increases.
+In Figure 3(b), we observe a similar pattern, though with a clearer picture on the distribution of false positives and negatives.
+Specifically, we see that the number of true positives is quite stable across many judges.
+The true negatives, instead, drop off quickly as the judge quality decreases, suggesting it is generally easier to judge answers that are correct.
+Figure 4: Precision, recall and prompt sensitivity.
+(a) Recall and precision improve with increasing human alignment ($R^{2}$ = 0.31 and $R^{2}$ = 0.21, respectively).
+(b) Scott’s $\mathbf{\pi}$ scores for judges across different instructions.
 Next, we analyse the types of errors made by the judge models by manually annotating 900 outputs from Llama-7B Base with error codes, focusing on the top performers GPT-4 Turbo and Llama-3 70B.
@@
 Under-specified and otherwise incorrect answers are most challenging for both judges, while answers with too few entities are judged relatively accurately by GPT-4 but less accurately by Llama-3 70B.
+Table 2: Error analysis for GPT-4 and Llama-3 70B judges.
+The example question is “Excluding Lady Jane Grey, who were the five monarchs of the House of Tudor?”, the correct answer “Henry VII, Henry VIII, Edward VI, Mary I and Elizabeth I” (in any order).
 Error code Explanation Example Proportion GPT-4 recall Llama-3 70B recall
@@
 1.23% 20.0% 40.0%
-Table 2: Error codes used to identify the types of errors made by exam-taker models when answering questions.
-The example question in this case is “Excluding Lady Jane Grey, who were the five monarchs of the House of Tudor?”, with the correct answer “Henry VII, Henry VIII, Edward VI, Mary I and Elizabeth I” (in any order).
 5.2 Judge model sensitivity to prompt length and specificity
-Next, we study the impact of the prompt on the ability of the judge models to perform an accurate assessment of exam-taker models, with two somewhat intertwined goals: 1) to study if the success of various judge models is related to the length of the prompt, and, 2) to study the degree to which the judgments of the judge models change with the specificity of the prompt.
+Next, we study the impact of the prompt on the predictions of the judge models, to understand if the success of various judge models is related to the length of the prompt, and to study the degree to which the judgments of the judge models change with the specificity of the prompt.
 We use four different prompt versions, varying in length and specificity.
-Each prompt instructs the judge models on how to evaluate responses, with complexity and detail increasing with the token count.
-The first two prompts (Without guidelines V1 and V2, with 45 and 58 tokens, respectively) simply ask to evaluate the responses, without any further information, while more elaborate guidance and examples are given in the longer prompts (Guidelines without examples and Guidelines with examples, with 245 and 301 tokens, respectively).
-All prompts are listed in Appendix K.
-Figure 4(b) shows that Mistral 7B (Jiang et al., 2023), GPT-4 Turbo and Llama-3 70B exhibit relatively low variance in their agreement with humans as the level of information and the length of the prompt increases.
-For this task, top performers’ (GPT-4 Turbo and Llama-3 70B) implicit definition of a correct judgment seems well aligned with the provided instructions and thus shows high alignment with humans even if no specific instructions are provided.
+The first two prompts (Without guidelines V1/V2, 45 and 58 tokens, respectively) simply ask to evaluate the responses, without any further information, while more elaborate guidance and examples are given in the longer prompts (Guidelines without examples and Guidelines with examples, 245 and 301 tokens, respectively).
+All prompts are listed in Appendix M.
+Figure 4(b) shows that GPT-4 Turbo, Llama-3 70B and Llama-3.1 70B exhibit relatively low variance in their agreement with humans as the level of information and the length of the prompt increases.
+For this task, top performers’ (GPT-4 Turbo, Llama-3 70B and Llama-3.1 70B) implicit definition of a correct judgment seems well aligned with the provided instructions and thus shows high alignment with humans even if no specific instructions are provided.
 It can also be observed that only top performers appears to benefit from the more detailed instructions, with a slight upward trend, whereas the other models get less aligned with more instructions.
 This might be due to the less powerful judges not being able to follow many instructions in the prompt at the same time.
-Interestingly, in a follow-up experiment (see Appendix L), when the order of references provided to the judge models is shuffled, Figure 6(b) demonstrates that larger judge models consistently maintain their judgments regardless of the reference order, whereas smaller models except Mistral 7B are more sensitive to the reference order given in the prompt.
+In a follow-up experiment, we further investigate the impact of the order of the reference answers (for details, we refer to Appendix N).
+Figure 6(b) illustrates that larger judge models consistently maintain their judgments regardless of the reference order, whereas smaller models – with the exception of Mistral 7B – are more sensitive to the reference order given in the prompt.
 5.3 Evaluating controlled responses
-Figure 5: We observe that judge models remain robust when exam-taker models produce responses identical to the prompt.
-However, this robustness diminishes in the presence of hallucinated responses such as "Yes" and "Sure" from the exam-taker models.
-Additionally, when assessing the gold standard answer, judges do not consistently arrive at the correct judgement 100% of the time.
-Next, we perform some simple tests on the judge models by asking them to evaluate a set of dummy benchmark responses.
-For the first test, the answer to be evaluated for each question is one of the references from the dataset, verbatinm (the answer is thus always correct), while for the next three tests, the answer to be evaluated is always incorrect, with the dummy exam-taker model always responding with ‘‘Yes’’, and ‘‘Sure’’ for the second and the third tests, respectively, and simply repeating the question for the fourth test.
-Figure 5 shows that while some judge models are able to identify and correctly mark the answers as correct (for the first test) or incorrect (for the next three tests), some judges, notably Llama-2 70B incorrectly evaluates a significant number of dummy answers, even though it shows a relatively high alignment with humans on the benchmark evaluations (see Figure 1(b)).
+Figure 5: Judge responses to dummy answers.
+We investigate how judge models respond to dummy answers.
+judge models remain robust when exam-taker models produce responses identical to the prompt (‘repeater’), but are less robust when the responses are "Yes" and "Sure".
+Even when the answer matches one of the reference answers verbatim (‘Gold answer’), judges do not always arrive at the correct judgement.
+Next, we perform simple tests on the judge models by asking them to evaluate a set of dummy benchmark responses.
+For the first test, the answer to be evaluated for each question is one of the references from the dataset, verbatim (the answer is thus always correct).
+For the next three tests, the answer is always incorrect.
+In the second and third tests, the dummy exam-taker model always responds with ‘‘Yes’’, and ‘‘Sure’’ for the second and third tests, respectively.
+For the fourth test, the evaluated answer is a repetition of the question.
+In Figure 5, we can see that while some judge models are able to identify and correctly mark the answers as correct (for the first test) or incorrect (for the next three tests), some judges, notably Llama-2 70B, incorrectly evaluate a significant number of dummy answers, even though they show a relatively high alignment with humans on the benchmark evaluations (see Figure 1(b)).
 We hypothesise that when the answers are plausible but incorrect (e.g.
@@
 However, the judges might get confused about what they are supposed to evaluate if the answer is completely unrelated to the question (such as the words ‘‘Yes’’ and ‘‘Sure’’).
-It is possible that in this situation a judge model tries to evaluate one of the reference answers, thus marking it as correct, though further research is required to identify the cause of this behavior.
+It is possible that, in this situation, a judge model tries to evaluate one of the reference answers, thus marking it as correct, though further research is required to identify the cause of this behavior.
 5.4 Leniency bias in judge models
-Judge model $\kappa$ $P_{c}$ $P_{+}$
-Gemma 2B 0.50 0.62 0.80
-Llama-2 7B 0.66 0.68 0.36
-Mistral 7B 0.72 0.75 0.75
-Llama-2 13B 0.68 0.76 0.45
-Llama-2 70B 0.80 0.79 0.94
-Llama-3 8B 0.81 0.80 0.84
-Llama-3 70B 0.84 0.84 0.79
-GPT-4 Turbo 0.85 0.85 0.66
-Figure 6: (a) Leniency bias across various judge models.
-To compute the leniency bias in judge models we estimate $P_{c}$ and $P_{+}$ for different judge models; (b) Consistency scores of each judge model for 3 random permutations of references in the prompt.
-The consistency score is the percentage of questions for which the judge model gives the same judgment for all 3 runs.
-We observe that the bigger models are more consistent as judges and less sensitive to reference orders in the prompt.
-More details are given in Appendix L.
-To get a general sense of the inherent biases or misalignment in the evaluation criteria that might be present in the judge models, we estimate if they have a positive or negative bias in their judgment.
+Judge model $\pi$ $P_{c}$ $P_{+}$
+JudgeLM-7B 0.65 0.68 0.19
+Mistral-7B 0.66 0.70 0.87
+Llama2-70B 0.69 0.66 0.99
+Llama2-13B 0.74 0.74 0.87
+Llama3.1-8B 0.77 0.77 0.82
+GPT-4 0.87 0.87 0.69
+Llama3.1-70B 0.88 0.88 0.82
+Llama3-70B 0.88 0.87 0.90
+Figure 6: Leniency bias and answer consistency.
+(a) Leniency bias across various judge models.
+(b) Consistency score, defined as the percentage of questions for which the judge model gives the same judgment for three different answer orders.
+Lastly, to get a general sense of the inherent biases or misalignment in the evaluation criteria that might be present in the judge models, we estimate if they have a positive or negative bias in their judgment.
 To do so, we assume that a judge assigns the correct judgment (i.e.
 same evaluation as the ground truth) with a probability of $P_{c}$ and assigns the rest of the samples to be ‘‘correct’’ with a probability $P_{+}$, which we call their leniency bias.
-We estimate the values of $P_{c}$ and $P_{+}$ from the benchmark results,444 The theoretical derivation of the expressions for $P_{c}$ and $P_{+}$, as well as the empirical validation for their estimated values can be found in Appendix M.
+We estimate the values of $P_{c}$ and $P_{+}$ from the benchmark results555 The theoretical derivation of the expressions for $P_{c}$ and $P_{+}$, as well as the empirical validation for their estimated values can be found in Appendix O.
 and show them in Figure 6(a).
@@
 6 Conclusion
-In this work, we provide an extensive study of the properties of LLMs as judges, comparing them with human evaluation and automated evaluation methods.
-By focusing on a relatively ‘clean’ evaluation scenario in which inter-human agreement is high, we examine the potential issues with the LLM-as-a-judge paradigm separately from the ambiguity and subjectivity in the task itself.
-We find that even in relatively straightforward setups, smaller and more cost-efficient models are less effective judges compared to the best available LLMs – also if they are specifically trained as judges, such as JudgeLM – and that even the best LLMs fail to meet the consistency of humans.
-Furthermore, though previous work commonly used percent agreement, we found that Cohen’s kappa ($\mathbf{\kappa}$) distinguishes judges much better.
-However, we observe that even judges with excellent $\mathbf{\kappa}$ scores may require further scrutiny.
-While GPT-4 Turbo and Llama-3 both have excellent alignment scores, simpler and more cost-efficient approaches like JudgeLM and contains perform better when discriminating between the exam-taker models in terms of their ranking, despite having much lower alignment scores and more systematic biases.
-In further analyses, we find that: 1) LLMs tend to judge positively when in doubt, and this is more pronounced for small models than for larger ones; 2) Judge models with lower alignment lack precision rather than recall; 3) GPT-4 Turbo and Llama-3 70B are generally robust across different prompts, but are difficult to ‘steer’ in their judgments; 4) Some judge models can be easily fooled by dummy answers such as ‘Yes’ and ‘Sure’; 5) Judge models are better at detecting completely incorrect answers than partially incorrect ones.
+In this work, we provide an extensive study of the properties of LLMs as judges, comparing them with human judges as well as automated evaluation methods.
+By focusing on a clean evaluation scenario in which inter-human agreement is high, we examine the potential issues with the LLM-as-a-judge paradigm, separately from the ambiguity and subjectivity in the task itself.
+We find that even in straightforward setups, smaller and more cost-efficient models are less effective judges compared to the best available LLMs, such as Mistral 7B.
+GPT-4 Turbo, Llama-3.1 70B and Llama-3  70B, instead, are much better aligned, though they are still quite far from the alignment that humans have among each other.
+In some cases, despite their high alignment, their scores deviate from human scores with up to 5 points.
+Given the relative simplicity of the scenarios in which we deployed the judges, urging caution in using judge for more complex scenarios.
+Importantly, we noted that such patterns are virtually undetectable using the commonly deployed metric of percent aligned, which barely discrimates between the considered judges.
+We suggest that future work instead considers the more robust metric Scott’s $\mathbf{\pi}$, which allows to distinguish judges much better.
+Next, we note that to discriminate between models, high alignment scores are not an absolute necessity.
+While GPT-4 Turbo and Llama-3 both have excellent alignment scores, simpler and more cost-efficient and even the lexical matching method contains perform on par when discriminating between the exam-taker models in terms of their ranking, despite having much lower alignment scores and score deviations.
+If the purpose of a study is to determine which model is better and not to estimate their actual scores, such approaches may thus be as suitable as the more expensive ones.
+Lastly, we run a range of experiments to investigate judge models’ sensitivity to prompts, their precision and recall, their error types, how lenient they are, and how much they can be fooled by dummy answers.
+We find that LLMs tend to judge positively when in doubt, and this is more pronounced for small models than for larger ones; that judge models with lower alignment lack precision rather than recall, that better models are generally more robust across different prompts, but are difficult to ‘steer’ in their judgments; that some judge models can be easily fooled by dummy answers such as ‘Yes’ and ‘Sure’; and that judge models are better at detecting completely incorrect answers than partially incorrect ones.
 Overall, this work adds to the realm of LLM evaluation research by assessing judges within a clearly defined and objective framework.
 Our results highlight the utility of using some LLMs as judges but also urge caution in blindly trusting their judgments, even if they are found to be well-aligned with humans.
-For practitioners using LLMs as judges – regardless of the setup – we recommend not only computing percent agreement, but also Cohen’s kappa, and pairing these with a qualitative analysis to ensure that conclusions from judge models are less susceptible to biases.
+For practitioners using LLMs as judges – regardless of the setup – we recommend not only computing percent agreement, but also Scott’s $\mathbf{\pi}$, and pairing these with a qualitative analysis to ensure that conclusions from judge models are less susceptible to biases.
 We further elaborate on the limitations of our work in Appendix A.
-In the future, we plan to expand our work to increasingly more complex scenarios with more open-ended answers and variability, and more generally assess how consistent our findings are across dataset samples, benchmarks, and prompt templates.
+In the future, we plan to expand our work to increasingly more complex scenarios with more open-ended answers and variability and more generally assess how consistent our findings are across dataset samples, benchmarks, and prompt templates.
 Appendix A Limitations
-In our work, we have shown potential pitfalls of using LLMs as judges, as well as the issues with commonly used metrics for quantifying their judgment quality.
-We have observed several trends and interesting phenomena regarding the behaviour of judge models in the LLM-as-a-judge paradigm, such as poor correlation or precision with alignment (§ 5.1), inability of smaller judge models to follow complex and nuanced instructions (§ 5.2), and tendencies of judge models to be lenient in their judgment when they are unsure (§ 5.4).
-Broadly, our results show that caution and scrutiny are required in using LLMs as judges, but not all our specific conclusions about rankings may hold across different examples, different judge models and exam-taker model accuracies, and tasks.
-For example, on a small experiment with a different sample and different instructions, we observed that when we limited the questions to only those that have fewer than 10 or 20 references, there was a consistent 5-10 point drop in percent-alignment across all judge models, but the discriminative ability of the top-performing judges improved.
-In that example, in which the overall accuracy of the models was substantially lower than in the example in our main experiment, we also observed a closer match between Cohen’s kappa and percent alignment, suggesting that the accuracies of the exam-taker models being compared may play an important role in how well their scores match.
+In our work, we have evaluated how 11 different LLMs fare as judges in a scenario in which judgements should be relatively straight-forward, and human alignment is high.
+As any study, our work has several limitations as well as directions that we did not explore but would have been interesting too.
+In this section, we discuss both.
+Simplicity of the task
+As mentioned in the introduction of our work, the scenario in which judges are used are typically much more complicated than the scenario that we focussed on.
+Specifically, judges are most often deployed in preference rankings (where two model responses are compared) or to judge complex answers that are difficult to automatically parse.
+In such tasks, human agreement is often low, making it challenging to judge the judges themselves.
+In our work, we have deliberately chosen for a simple task, in which human alignment is high.
+The main premise is, that if a judge does not perform well in this simple setup, caution is suggested also in more complex setups – if someone cannot do multiplication, why would they be able to solve ordinary differential equations.
+Given the poor understanding of which abilities of LLMs generalise in what dimensions, however, more studies are needed to understand how our results generalise to various other scenarios.
+Human alignment
+In an earlier version of this paper, due to the high cost of human annotations, we opted to select a single model for human annotation as we iteratively modified the exam taker prompt, few-shot examples, and guidelines.
+We selected the Llama2 7B for this purpose with a random sample of 600 questions.
+As this is only a single model, it is possible that our human alignment scores are biased because of that.
+After, we have therefore extended our results with another 600 human-annotated examples from Llama3.1 70B.
+For Llama2 7B The average alignment among human evaluators had a Scott’s $\mathbf{\pi}$ of $96.36\pm 1.46$,and the average percent agreement was $98.33\%\pm 0.76\%$.
+For Llama3.1 70B, we noted that the average alignment among human evaluators had Scott’s $\mathbf{\pi}$ of $95.78\pm 0.30$,and the average percent agreement was $98.72\%\pm 0.10\%$.
+Given the similarity of these two numbers, we believe that these 1200 samples provide an adequate estimate.
+In the paper, we take the average.
+Size of the judged samples
+As each of the nine exam-taker models requires human annotations for each sample, we restricted our analysis to 400 samples in total.
+This sample size also allowed us to conduct manual annotations and error analysis within 75 human hours/200 GPU hours (see Appendix H) and give reliable confidence intervals while also providing the flexibility to compare a range of models.
+We were not able to increase the size due to the cost, but a statistical analysis (details provided in Appendix I) illustrated that the variance because of this sample size was very low.
+Selection of judges
+With our selection of judges, we have stuck to autoregressive judges that can be used off-the-shelve, as well as one LLM specifically trained to judge.
+They are – at the moment of writing – the ones that are most commonly used as LLM-judges, and we have tried to be comprehensive across size and family.
+Nevertheless, we acknowledge that there are other judges that we could have considered as well.
+As including more judges in – compared to including more exam-taker models– relatively straightforward because it requires only computational power, no manual annotation, we hope that others may evaluate their newly proposed judges using our setup as well.
+Future work
 All in all, these differences underline how finicky using LLMs as judges can be, and with that confirm the overall conclusions of our study that much more work is needed to better understand the strengths and limitations of judge models across a wide range of scenarios and model accuracies.
 We consider assessing the strengths across multiple different samples and tasks, which would require many more human annotations, outside the scope of this paper and leave such experimentation for future work.
-Appendix B Model and dataset details
+Appendix B A brief explanation of the theoretical issues with Cohen’s kappa
+Cohen’s Kappa Coefficient (Cohen, 1960) is a statistic to measure inter-rater agreement for categorical responses.
+Cohen’s Kappa coefficient measures this agreement by computing the observed (percent) agreement between raters ($p_{o}$) and comparing it with the hypothetical probability of chance agreement ($p_{e}$), which is taken as a baseline, as follows:
+$\kappa\equiv\frac{p_{o}-p_{e}}{1-p_{e}}$ (1)
+In this equation, the chance agremeent $p_{o}$ constitutes the hypothetical probability that observed agreement occurred by chance, given the observed distributions of the considered raters, under the assumption that the probabilities the raters assign to the observed labels are independent.
+Specifically, it is defined as:
+$p_{e}=\sum_{k}\widehat{p_{k12}}=^{ind}\sum_{k}\widehat{p_{k1}}\widehat{p_{k2}}=\sum_{k}\frac{n_{k1}}{N}\frac{n_{k2}}{N}=\frac{1}{N^{2}}\sum_{k}n_{k1}n_{k2},$ (2)
+where $\widehat{p_{k12}}$ is the estimated probability that rater 1 and rater 2 will classify the same item as $k$, rewritten to $\widehat{p_{k1}}\widehat{p_{k2}}$ under the assumption that $p_{k1}$ and $p_{k2}$ are independent.
+The crux of the issue with this method of computation, is that $\widehat{p_{k1}}$ and $\widehat{p_{k2}}$ are estimated independently from the data.
+As such, the chance agreement adjusts for the observed average differences between raters, which is in fact part of what we intend to measure.
+To address this issue, Scott’s Pi (Scott, 1955) instead defines the chance baseline under the assumption that the raters have the same distribution, which is estimated considering the joint distribution of rater 1 and rater 2, rather than considering them separately.
+It defines $p_{e}$ as:
+$p_{e}=\sum_{k}\widehat{p_{k}^{2}}=\sum_{k}\sum_{k}(\frac{n_{k1}+n_{k2}}{2N})^{2}$ (3)
+As such, contrary to Cohen’s Kappa, it captures differences surpassing the chance agreement if rater 1 and rater 2 were in fact equivalent.
+In other words, we compare against a baseline in which raters would be equivalent, and we measure how much they deviate from that.
+Note that if the empirical distributions of rater 1 and rater 2 are the same, so will the values of Scott’s Pi and Cohen’s Kappa be.
+This also implies that for larger observed (percent) alignment values, the values for Cohen’s Kappa and Scott’s Pi will be closer.
+Appendix C Model and dataset details
 In Table 3, we show the different models and datasets used in our experiments, along with version and license details.
@@
 Llama-3 70B Chat meta-llama/Meta-Llama-3-70B-Instruct llama3
+Llama-3.1 8B Chat meta-llama/Meta-Llama-3.1-8B-Instruct llama3.1
+Llama-3.1 70B Chat meta-llama/Meta-Llama-3.1-70B-Instruct llama3.1
 JudgeLM BAAI/JudgeLM-7B-v1.0 Non-commercial license
 GPT-4 Turbo gpt-4-turbo-2024-04-09 N/A
-Qwen 0.5B Chat Qwen/Qwen1.5-0.5B-Chat tongyi-qianwen
-Qwen 1.8B Chat Qwen/Qwen1.5-1.8B-Chat tongyi-qianwen
-Qwen 4B Chat Qwen/Qwen1.5-4B-Chat tongyi-qianwen
-Qwen 7B Chat Qwen/Qwen1.5-7B-Chat tongyi-qianwen
-Qwen 14B Chat Qwen/Qwen1.5-14B-Chat tongyi-qianwen
-Qwen 72B Chat Qwen/Qwen1.5-72B-Chat tongyi-qianwen
-Appendix C Model evaluation prompt templates
+Appendix D Model evaluation prompt templates
 In Figure 7 and Figure 8, we show the prompt templates used for the base and chat exam-taker models during the question answering process.
@@
 Figure 8: Prompt template for Chat exam-taker models
-Appendix D Judge LLM Prompt templates
+Appendix E Judge LLM Prompt templates
 In Figure 9, we show the prompt template used to guide the judge models during the evaluation process of a 400-question sample from the TriviaQA unfiltered dataset.
 Figure 9: Prompt templates for the judge models
-Appendix E Metrics for judge models
+Appendix F Metrics for judge models
 If one of the annotators is taken to be the reference, then the annotations of the other annotator can be categorized as true positives, false positives, true negatives, and false negatives, with the total number of each of them in a benchmark being represented by $T_{P},F_{P},T_{N},$ and $F_{N}$ respectively.
@@
 $\rho=\frac{T_{P}+T_{N}}{T_{P}+F_{P}+T_{N}+F_{N}}.$ (4)
-Cohen’s kappa coefficient, or Cohen’s kappa for short (Cohen, 1960), measures the alignment of two annotators while also taking into account the possibility of agreement by pure chance.
-This coefficient can have values between $-1$ and $1$, but is usually above 00 in most real-world situations.
-The value of Cohen’s kappa is given as
-$\kappa$ $=\frac{p_{o}-p_{e}}{1-p_{e}}=\frac{2(T_{P}T_{N}-F_{P}F_{N})}{(T_{P}+F_{P})(T_{N}+F_{P})+(T_{P}+F_{N})(T_{N}+F_{N})}.$ (2)
+Scott’s Pi, (Scott, 1955), measures the alignment of two annotators while also taking into account the possibility of agreement by pure chance.
+This coefficient usually has values above 00 in most real-world situations.
+The value of Scott’s Pi is given as
+$\pi$ $=\frac{p_{o}-p_{e}}{1-p_{e}}=\frac{\left(\frac{TP+TN}{TP+TN+FP+FN}\right)-\left(\frac{(TP+FP)(TP+FN)+(TN+FN)(TN+FP)}{(TP+TN+FP+FN)^{2}}\right)}{1-\left(\frac{(TP+FP)(TP+FN)+(TN+FN)(TN+FP)}{(TP+TN+FP+FN)^{2}}\right)}$ (5)
 Here, $p_{o}$ is the relative observed agreement, and $p_{e}$ is the hypothetical probability of chance agreement.
-This coefficient is considered to be a more robust measure of inter-annotator alignment, but also less interpretable in terms of what a particular value of $\kappa$ means.
-Generally, values of $\kappa$ in ranges $[0,0.2)$, $[0.2,0.4)$, $[0.4,0.6)$, $[0.6,0.8)$, and $[0.8,1)$ are considered to indicate no alignment, slight alignment, moderate alignment, substantial alignment, and near-perfect alignment respectively, with $\kappa=1$ indicating perfect alignment.
-Appendix F Human Annotation Guidelines
+Appendix G Human Annotation Guidelines
 We provide the guidelines used for human evaluation below.
@@
 If you have trouble judging whether the answer is correct, for instance because you feel you are lacking knowledge required to judge so, please indicate so by marking the answer "maybe correct" or ‘‘maybe incorrect", so that we can further review it.
-Appendix G Experiment costs
+Appendix H Experiment costs
 The costs for the different experiments described in this work belong in three categories – GPU-hours for running open-source models on one or more Nvidia A100 GPUs, OpenAI credits for making API calls to OpenAI models, [^6] Pricing details for OpenAI models are available at https://openai.com/api/pricing/ and human hours for manual annotations of benchmark responses.
@@
 Main evaluations 30 8 10
-Human alignment 1 - 6
+Human alignment 2 - 9
 Error analysis 1.5 - 5
@@
 Reference bias 5 4 1
-Total 72.5 24 23
-Appendix H Judge Scores
+Total 73.5 24 26
+Appendix I Statistical reliability of Evaluation sample
+Due to computational constraints discussed in Appendix A and Appendix H, we limit our evaluation set to randomly sampled 400 questions from TriviaQA (Joshi et al., 2017).
+In this section, we further take 5 samples of 300 randomly selected questions from the evaluation set and calculate the mean and standard deviation of Scott’s Pi.
+From Table 5, it can be observed that even on down-sampled sets, the Scott’s $\mathbf{\pi}$ values are similar to Figure 1(b).
+Standard deviation of all the judge models from the mean Scott’s $\mathbf{\pi}$ is also minimal, barring EM lexical match.
+Table 5: Weak Scott’s $\mathbf{\pi}$ variation for the 5 down-sampled sets indicating robustness for the evaluation sample
+Judge Model Mean Scott’s $\mathbf{\pi}$ Std Dev
+Llama3-70B 0.88 0.0046
+Llama3.1-70B 0.88 0.0039
+Llama3.1-8B 0.78 0.0050
+Llama2-13B 0.75 0.0043
+Llama2-70B 0.69 0.0114
+Mistral-7B 0.67 0.0108
+JudgeLM-7B 0.66 0.0026
+Contains 0.64 0.0087
+Llama3-8B 0.60 0.0126
+Llama2-7B 0.47 0.0112
+EM 0.47 0.29
+Gemma-2B 0.26 0.007
+Appendix J Judge Scores
 We show the scores assigned by each judge model to each exam-taker model, visualised in Figure 1(a) in Table 6.
@@
 Judge Models 7B 13B 70B 7B 13B 70B 7B
-Llama 3 8B 68.75 76.00 85.25 52.91 75.0 78.5 62.5 66.00 76.25
-Llama 3 70B 66.00 76.50 87.00 61.52 68.50 78.25 73.5 65.00 94.50
-Llama 2 7B 61.75 71.25 80.25 58.48 77.25 58.00 68.25 80.50 86.75
-Llama 2 13B 53.00 61.50 66.50 46.07 27.75 39.50 57.50 39.00 73.50
-Llama 2 70B 68.25 76.30 86.50 64.05 80.04 82.25 77.00 74.25 95.25
-Mistral 7B 67.75 75.00 85.25 64.55 68.75 77.50 74.50 68.50 92.75
-Gemma 2B 67.00 74.00 84.00 100.00 44.25 75.25 100.00 79.50 99.75
-JudgeLM 52.50 60.50 64.00 41.52 32.50 60.00 57.25 45.5 68.50
-GPT-4 63.25 75.00 85.0 56.7 57.50 72.0 72.5 52.75 92.25
-Exact Match 46.75 56.00 63.75 24.05 0.25 36.25 59.50 20.25 58.25
-Contains Match 50.75 60.00 68.00 38.98 46.25 59.50 57.25 44.00 70.00
-Human Eval 62.25 72.75 83.75 56.00 56.50 72.25 71.75 60.75 91.50
-Appendix I Exam-taker model base vs chat analysis
+Llama 3.1 8B 65.25 75.00 83.50 60.25 70.50 75.50 73.75 59.00 89.00
+Llama 3.1 70B 62.00 74.25 85.00 55.50 64.75 74.00 72.25 60.50 92.25
+Llama 3 8B 76.00 83.25 91.50 73.25 82.75 85.25 81.75 76.0 97.25
+Llama 3 70B 64.25 75.50 86.50 57.00 64.00 75.75 73.5 62.50 92.75
+Llama 2 7B 80.50 85.25 92.00 80.50 70.75 90.75 84.00 83.25 97.75
+Llama 2 13B 68.25 75.50 86.50 63.25 62.75 77.50 74.50 67.50 93.5
+Llama 2 70B 71.25 80.5 90.25 67.50 74.75 81.25 80.0 72.5 96.75
+Mistral 7B 72.50 80.75 90.50 69.00 74.75 82.50 80.25 72.00 96.25
+Gemma 2B 79.75 87.00 91.25 58.50 41 68.50 84.0 55.75 80.50
+JudgeLM 69.50 77.75 86.25 63.75 48.0 82.75 77.25 71.0 94.50
+GPT-4 60.50 71.50 82.50 54.50 59.0 73.0 69.75 56.50 90.0
+Exact Match 46.75 56.00 63.75 24.00 0.25 36.25 59.50 20.25 58.25
+Contains Match 50.75 60.00 68.00 39.00 46.25 59.50 57.25 44.00 70.00
+Human Eval 62.50 72.75 83.75 56.00 56.50 72.25 71.75 60.75 91.50
+Appendix K Exam-taker model base vs chat analysis
 Table 7: Scores of base and chat models by various judges
@@
 Base Chat Base Chat Base Chat Base Chat Base Chat
-Llama-2 7B 46.75 24.05 50.75 38.98 62.25 56.00 63.25 56.75 66.0 61.50
-Mistral 7B 59.50 20.25 57.25 44.00 71.75 60.75 72.50 52.75 73.5 65.0
-Llama-2 13B 56.00 0.25 60.00 46.25 72.75 56.50 75.00 57.50 76.50 68.50
-Llama-2 70B 63.75 36.25 68.00 59.50 83.75 72.25 85.00 72.00 87.00 78.25
+Llama-2 7B 46.75 24.00 50.75 39.00 62.25 56.00 60.50 54.50 64.25 57.00
+Mistral 7B 59.50 20.25 57.25 44.00 71.75 60.75 69.75 56.50 73.50 62.50
+Llama-2 13B 56.00 0.25 60.00 46.25 72.75 56.50 75.00 59.00 76.50 64.00
+Llama-2 70B 63.75 36.25 68.00 59.50 83.75 72.25 82.50 73.00 86.50 75.75
 Given the human judgments we have available, we take the opportunity to investigate the performance differences between base and their corresponding chat models.
@@
 Substantially more often than the base models, the chat models do answer the question with a semantically plausible but incorrect entity.
-In Table 7-9, we provide examples of such cases.
+In Table 8-Table 10, we provide examples of such cases.
 The results do not show any evidence to support the first hypothesis: the number of errors where the answer cannot be parsed or is just entirely incorrect does not differ between base and chat models.
@@
 Mistral 7B Chat Jack Walker & Sons
-Appendix J Exam-taker model ranking correlation
-In Table 10, we show the Spearman’s rank correlation coefficient (Spearman, 1904) ($\rho$) with human judgment.
-Since $\rho$ > 0.7 is considered well aligned, only Llama-7B and Gemma-2B have poor rank correlation with human judgment.
+Appendix L Exam-taker model ranking correlation
+In Table 11, We use the Spearman Rank correlation coefficient (Spearman, 1904) to assess the rankings of the exam-taker models.
+To validate these rankings, we randomly select 6 out of 9 exam-taker models across 5 samples, subsequently calculating the mean ($\rho$) and standard deviation ($\sigma$) of the rankings.
+The results reveal that the contains model exhibits the highest stability and $\rho$ among the rankings, while the majority of judge models achieve a coefficient exceeding 0.7, indicating a strong alignment.
+Notably, smaller models such as Mistral 7B perform on par with GPT-4 Turbo, highlighting the robustness of smaller models in maintaining rankings.
 Table 11: Spearman Rank Correlation Coefficient $\rho$.
-Judges $\rho$
-Contains 0.98
-JudgeLM-7B 0.98
-GPT-4 0.93
-Llama3-70B 0.93
-Mistral-7B 0.92
-Llama-13B 0.82
-EM 0.78
-Llama3-8B 0.77
-Llama-70B 0.75
-Llama-7B 0.39
-Gemma-2B 0.21
-Appendix K Too much info confuses judges
+Judges $\rho$ $\sigma$
+Contains 0.99 0.02
+Mistral-7B 0.98 0.03
+GPT-4 0.98 0.03
+Llama2-13B 0.95 0.18
+JudgeLM-7B 0.95 0.05
+Llama2-7B 0.94 0.04
+Llama3.1-70B 0.94 0.07
+Llama3-70B 0.93 0.05
+Llama3.1-8B 0.89 0.10
+Llama3-8B 0.86 0.07
+Llama2-70B 0.84 0.13
+Gemma-2B 0.71 0.20
+EM 0.67 0.13
+Appendix M Too much info confuses judges
 In Figure 11-14, we report the guidelines we used for the experiments in § 5.2.
@@
 Figure 14: Guidelines with Examples Prompt template for the judge models
-Appendix L Judge models are sensitive to reference order
+Appendix N Judge models are sensitive to reference order
 We investigate the judges’ sensitivity to reference order by providing the same prompt, question and model response to the judge models, but shuffling the reference order in three different permutations.
@@
 Figure 16: Example of Llama2-7B failing to identify the task by changing the order of the references.
-Appendix M Leniency Bias
+Appendix O Leniency Bias
 As described in § 5.4, for the purpose of the leniency bias experiments, we assume that a judge assigns the correct judgment with a probability of $P_{c}$ and randomly assigns the rest of the samples to be “correct” with a probability $P_{+}$.
@@
 The values of $P_{c}$ and $P_{+}$ can be estimated from observed data using the derived expressions.
-In this experiment, we include Qwen models (Bai et al., 2023) of varying sizes, in our judge ensemble to increase the number of data points for this study.
 The estimated probabilities using this method, with human evaluation as the reference, are shown in Figure 17(a).
-To validate these derived values, we observe the correlation between the estimated values of $P_{c}$ and Cohen’s kappa ($\kappa$).
-As shown in Figure 17(b), we observe that the estimated values of $P_{c}$ are highly correlated to the Cohen’s kappa values for the judge models, with a Pearson correlation coefficient of $0.98$.
-Judge model $\kappa$ $P_{c}$ $P_{+}$
-Qwen 1.8B 0.19 0.18 1.00
-Qwen 0.5B 0.19 0.19 0.87
-Gemma 7B 0.40 0.32 0.07
-Gemma 2B 0.50 0.62 0.80
-Qwen 4B 0.69 0.67 0.89
-Llama-2 7B 0.66 0.68 0.36
-Mistral 7B 0.72 0.75 0.75
-Llama-2 13B 0.68 0.76 0.45
-Llama-2 70B 0.80 0.79 0.94
-Qwen 7B 0.80 0.79 0.77
-Llama-3 8B 0.81 0.80 0.84
-Qwen 72B 0.82 0.80 0.93
-Qwen 14B 0.83 0.82 0.83
-Llama-3 70B 0.84 0.84 0.79
-GPT-4 Turbo 0.85 0.85 0.66
-Figure 17: a) Estimated values of $P_{c}$ and $P_{+}$ for different judge models. b) Pearson’s correlation coefficient between $\kappa$ and $P_{c}$ for judge models.
+To validate these derived values, we observe the correlation between the estimated values of $P_{c}$ and Scott’s Pi ($\pi$).
+As shown in Figure 17(b), we observe that the estimated values of $P_{c}$ are highly correlated to the Scott’s $\mathbf{\pi}$ values for the judge models, with a Pearson correlation coefficient of $0.98$.
+Judge model $\pi$ $P_{c}$ $P_{+}$
+Gemma-2B 0.26 0.38 0.87
+Llama2-7B 0.47 0.63 0.75
+Llama3-8B 0.59 0.63 0.74
+JudgeLM-7B 0.65 0.68 0.19
+Mistral-7B 0.66 0.70 0.87
+Llama2-70B 0.69 0.66 0.99
+Llama2-13B 0.74 0.74 0.87
+Llama3.1-8B 0.77 0.77 0.82
+GPT-4 0.87 0.87 0.69
+Llama3.1-70B 0.88 0.88 0.82
+Llama3-70B 0.88 0.87 0.90
+Figure 17: a) Estimated values of $P_{c}$ and $P_{+}$ for different judge models. b) Pearson’s correlation coefficient between $\pi$ and $P_{c}$ for judge models.